### v1

# Market-Driven Equilibria for Distributed Solar Panel Investment

Abstract
This study investigates market-driven long-term investment decisions in distributed solar panels by individual investors. We consider a setting where investment decisions are driven by expected revenue from participating in short-term electricity markets over the panel’s lifespan. These revenues depend on short-term markets equilibria, i.e., prices and allocations, which are influenced by aggregate invested panel capacity participating in the markets. We model the interactions among investors by a non-atomic game and develop a framework that links short-term markets equilibria to the resulting long-term investment equilibrium. Then, within this framework, we analyze three market mechanisms: (a) a single-product real-time energy market, (b) a product-differentiated real-time energy market that treats solar energy and grid energy as different products, and (c) a contract-based panel market that trades claims/rights to the production of certain panel capacity ex-ante, rather than the realized solar production ex-post. For each, we derive expressions for short-term equilibria and the associated expected revenues, and analytically characterize the corresponding long-term Nash equilibrium aggregate capacity. We compare the solutions of these characterizing equations under different conditions and theoretically establish that the product-differentiated market always supports socially optimal investment, while the single-product market consistently results in under-investment. We also establish that the contract-based market leads to over-investment when the extra valuations of users for solar energy are small. Finally, we validate our theoretical findings through numerical experiments.
I Introduction
The increasing adoption of renewable energy, particularly solar photovoltaic systems, plays an important role in modern power systems. While utility-scale solar projects contribute significantly to overall capacity, distributed rooftop installations are playing an increasingly prominent role in reshaping the electricity landscape. Notably, as of 2022, approximately 140 GW of solar capacity had been installed in the U.S., with more than 35% of it coming from installations by households and commercial entities [1].
Unlike centralized projects planned and operated by utilities, distributed solar investments are driven by decisions of individual households and businesses. Several factors influence these decisions. A key factor is the financial return from the investment, driven by the expected revenue earned in short-term distribution-level markets over the panels’ lifespan, where that revenue is determined by the markets’ equilibria. Meanwhile, these short-term equilibria depend on the aggregate invested panel capacity available in the markets, which is shaped by the long-term investment equilibrium that emerges from interactions among individual investors. Therefore, there is a feedback loop between short-term markets equilibria and long-term investment equilibrium, as each influences the other. A precise characterization of this relationship is crucial for understanding the emergence of investment equilibria. This insight is not only important for utilities and independent system operators, who should account for distributed capacity growth in their planning studies, but also for policymakers and market designers aiming to understand the long-term impact of diverse market designs on the solar panel adoption. This motivates us to study the following critical open question: How do short-term market equilibria, arising from diverse solar market mechanisms, shape the corresponding long-term distributed investment equilibrium?
I-A Organization and Contributions
To address this problem, we consider a large population of individual investors and aim to characterize the long-term investment equilibrium that emerges from the previously described feedback loop. We first develop an investment model, formulated as a non-atomic game, that links long-term investment decisions to expected revenues under a given market design (Section II). To characterize these revenues, we introduce a unified modeling framework for short-term markets (Section III). This framework allows us to formally define and compare different market designs by capturing key features such as product differentiation and ex-ante/ex-post trading. We then apply it to three representative mechanisms that differ in these key features (Section IV): (a) a single-product real-time market, (b) a product-differentiated real-time market, and (c) a contract-based panel market. For each, we first derive analytical characterizations of the short-term market equilibria and then use the investment model to derive the associated long-term equilibrium aggregate capacities (Section V). Our analysis, initially assuming homogeneous operation periods, is extended to handle heterogeneous operation periods (Section VI), followed by numerical experiments (Section VII).
We contribute to the literature in the following key ways:
  1. (a)
    The non-atomic game-based investment model encapsulates how individual investors impact each other through their aggregate investment decisions, offering an analytically tractable way to link short-term market equilibria to long-term aggregate investment equilibrium.
  2. (b)
    The proposed unified modeling structure provides closed-form characterizations of short-term market equilibria across diverse mechanism designs, accommodating both ex-ante and ex-post trading as well as heterogeneous consumer preferences.
  3. (c)
    These analytical characterizations enable us to theoretically establish that the single-product real-time market, where solar energy is pooled with conventional grid energy and traded ex-post, leads to under-investment. In contrast, the product-differentiated market, where solar energy is traded ex-post and may be sold at a premium due to heterogeneous user preferences, yields the socially optimal capacity. Finally, the contract-based market, where ex-ante panel capacity is traded instead of realized generation, can result in over-investment when users’ extra valuations for solar are low.
I-B Related Literature
This work is related to the intersection of two broad lines of literature: distribution-level electricity market mechanisms and long-term investment in solar panels. Most existing studies on the investment side of the literature adopt a centralized perspective, focusing on the optimal strategy of a single decision-maker such as a utility planner or an individual investor (see, e.g., [2, 3, 4, 5, 6]). However, this line of work neglects the interdependencies of investment decisions in distributed settings, where the actions of one investor can influence the returns of others through shared market mechanisms.
On the market side, a variety of distribution-level electricity market designs have been proposed. Understanding how these structures differ is essential for our analysis, as their design features shape the expected revenue of investors. In this context, prior research has proposed a range of market designs, with particular emphasis on real-time pricing and service contract-based approaches. In a broad sense, real-time pricing mechanisms involve the determination of time-varying electricity prices based on the ex-post realized values of supply and demand. In contrast, service contract-based mechanisms typically offer less volatile prices in exchange for a fixed level of electricity service, relying on ex-ante statistics such as those related to random solar generation. Extensive research has explored the relative merits of these retail market designs, examining factors such as economic efficiency [7, 8, 9, 10], price and bill volatility [11, 12, 13], and environmental impacts [14]. In addition, some studies have examined another facet of market design: whether electricity from solar and conventional sources should be treated as differentiated products in electricity markets, given their distinct operational characteristics and emissions [15, 16, 17, 18, 19]. Despite the contributions of all these reviewed studies, they all focus primarily on short-term performance criteria and overlook the long-term impacts of their designed markets on investment decisions.
The most closely related paper to ours is [20], which analyzes the equilibrium panel investment capacities under feed-in-tariff and sharing economy markets. Nonetheless, they adopt a finite-player game to analyze the setup involving large investors. In contrast, we focus on the setting with a large number of small investors. Through a non-atomic game formulation, we can analytically compare equilibrium capacities driven by various market mechanisms, for which they have to resort to numerical simulation.
This paper generalizes our previous conference work [21] by (a) extending our theoretical results to setups with heterogeneous operation periods, (b) presenting rigorous mathematical proofs, and (c) demonstrating our theoretical results by performing numerical experiments.
II Investment Model
A large group of potential investors considers installing solar panels and becoming panel owners. Their decisions weigh upfront capital and installation costs against expected revenues from participating in short-term electricity markets over the panels’ lifespan. An overview of the setup is shown in Fig. 1.
Figure 1: Schematic for distributed panel investment
II-A Connecting Planning and Operation Timescales
Investors typically consider a 25-year lifespan for solar panels [22]. However, estimating expected revenue at the time of investment is challenging, as reliable long-term forecasts of solar output and market conditions are rarely available. A common approach is to use a shorter planning period (e.g., one year), estimate revenues from historical data, and scale [^1] This step may include interest rate adjustments and net present value computations. the result by the number of such intervals over the panel’s lifespan. The appropriate length of the planning period should be selected based on the availability and granularity of historical data.
When the planning period is long, there may be multiple operation periods within it. We consider a finite-horizon discrete-time planning window consisting of operation periods $\mathcal{T}=\{1,\dots,T\}$, indexed by $t$, where $T$ is the total number of such periods. For each operation period $t\in\mathcal{T}$, we model the short-term operation conditions (e.g., supply and demand) as constants and incorporate variability (e.g., of solar generation) across different operation periods by sampling from a probability distribution. Thus, the expected revenue remains consistent across operation periods. We estimate it for a single representative period (e.g., as shown in Fig. 1), then scale it up to the planning horizon and ultimately to the full lifespan. The total expected revenue is therefore obtained by multiplying the representative period’s revenue by a scaling factor $\widetilde{T}$, which captures the number of such periods over the panel’s lifespan and incorporates any interest-rate adjustments.
Remark 1 (More granular operation model)
A more detailed operational model can be implemented by incorporating multiple representative periods, each characterized by distinct supply and demand values/distributions. For instance, separate solar generation profiles can be used for daytime and nighttime hours, along with different load patterns for weekdays and weekends. These settings can be accommodated within our framework; see Section VI for further discussion.
II-B Panel Investment Game
Considering a large number of small investors each of whom does not have market power, we model the group of investors as the continuous interval $\mathcal{I}_{\mathrm{inv}}:=[0,1]$. We consider the case where each investor $i\in\mathcal{I}_{\mathrm{inv}}$ decides whether to install a solar panel with a fixed capacity $\overline{c}$. Let the decision of investor $i$ be $x_{i}\in\{0,1\}$. The total panel capacity that the investors will install is then
$c=\overline{c}\int_{\mathcal{I}_{\mathrm{inv}}}x_{i}\,\,\mathrm{d}i.$ (1)
Let the capital and installation cost for a unit panel capacity be $\pi_{0}$. We assume that the panel upfront cost is attractive, at least when its solar production can be sold at the utility rate:
${\color[rgb]{0,0,0}c}\pi_{0}\leq\widetilde{T}\pi_{\mathrm{u}}\mathbb{E}[{\color[rgb]{0,0,0}c}G],$ (2)
where $\pi_{\mathrm{u}}>0$ is the fixed price at which the utility provides backstop electricity to buyers, and $G$ denotes solar generation per unit panel capacity, typically measured in $\mathrm{W}/\mathrm{m}^{2}$ or $\mathrm{GW}/\mathrm{km}^{2}$. This solar generation is modeled as a random variable with a cumulative distribution function $F_{G}$ and a probability density function $f_{G}$. For simplicity, we assume that $f_{G}$ is differentiable and that $G$ is supported on $\mathbb{R}_{+}$ with finite first and second moments, and $f_{G}(g)>0$ for all $g\in\mathbb{R}_{+}$.
The expected return for each investor $i$ is influenced not only by their own decision, $x_{i}$, but also by the collective actions of other participants and the specific short-term market mechanism assumed for the representative operation period. Thus, the payoff for investor $i$ from installing panel capacity $\overline{c}$ and selling the generated electricity in a short-term market $m$ (where $m$ serves as an index for different short-term market mechanisms) is given by the expected total revenue minus the investment cost, expressed as:
$\Pi^{\mathrm{inv}}_{m,i}(x_{i},c)=\left[(\Pi_{m}^{\mathrm{s}}(c)/c)-\pi_{0}\right]\overline{c}\,x_{i},$ (3)
where $\Pi_{m}^{\mathrm{s}}(c)$ represents the expected total revenue received by solar panel owners under the short-term market $m$.
Given the players $\mathcal{I}_{\mathrm{inv}}$, actions $x=\{x_{i}\}_{i\in\mathcal{I}_{\mathrm{inv}}}$, and payoff (3), we have defined the solar panel investment game, which is an aggregate game where the dependence of investor $i$’s payoff on the other investors’ decisions is only through the total panel capacity $c$. A Nash equilibrium (NE) for this game is defined as:
Definition 1 (NE for panel investment game)
For a market mechanism $m$, the aggregate panel capacity $c$ constitutes an NE for the solar panel investment game with market structure $m$ if there exists a collection of investment decisions $x$ such that $\Pi_{m,i}^{\mathrm{inv}}(x_{i},c)\geq\Pi_{m,i}^{\mathrm{inv}}(x_{i}^{\prime},c)$, for all $x_{i}^{\prime}\in\{0,1\}$ and $i\in\mathcal{I}_{\mathrm{inv}}$, and (1) holds.
Under this condition, no investor will have an incentive to unilaterally deviate from the NE decision given the aggregate capacity that is consistent with everyone’s investment decision.
Building on this investment model, we next examine several representative solar market mechanisms to assess how their structural differences influence investors’ payoffs and, in turn, the resulting long-term NE aggregate capacities.
III Unified Modeling Framework for Short-term Market Mechanism Analysis
We present generic modeling elements applicable across diverse market settings and then provide a detailed description of the specific market mechanisms analyzed in this paper.
III-A Common Model Elements
Consider an operation period (e.g., the representative operation period in Fig. 1) where the expected revenue of investors is shaped by short-term conditions in the distribution-level solar energy market. The market consists of buyers (i.e., electricity consumers), sellers [^2] We acknowledge the possibility that electricity prosumers can be both buyers and sellers at the same time. For the market mechanisms that we will consider in this paper, a prosumer can be modeled as the superposition of a buyer and a seller. This observation may not apply to other markets where buying and selling prices are non-uniform. (i.e., panel owners), and a utility company that provides backstop supply when solar generation alone cannot fully satisfy demand. Similar to the investment game, we model the groups of a large number of buyers and sellers as continuous intervals.
Define the set of buyers by the interval $\mathcal{I}_{\mathrm{b}}:=[0,1]$. In our analysis, we consider a situation where environmentally conscious electricity consumers have an extra heterogeneous valuation for solar energy over the conventional energy supplied by the utility. For a buyer $i\in\mathcal{I}_{\mathrm{b}}$, we denote their extra valuation for each unit of load served by solar as $v_{i}\in\mathbb{R}_{+}$, referred to as the solar premium. The distribution of $v_{i}$ can be characterized by a function $F_{V}$ defined as $F_{V}(v)=\int_{i\in\mathcal{I}_{\mathrm{b}}}\,\mathbb{1}\{v_{i}\leq v\}\,\,\mathrm{d}i$, where $\mathbb{1}$ is the indicator function. For simplicity, we assume that $F_{V}$ is differentiable and strictly increasing on $0,\overline{v}]$, where $\overline{v}\geq 0$. We denote its derivative by $f_{V}$, supported on $0,\overline{v}]$, and assume $f_{V}(v)$ is bounded away from zero for all $v\in[0,\overline{v}]$. Let the total electric load of all consumers be $L$. We assume $L$ is inelastic, deterministic, and equally distributed across all consumers.
Let the set of sellers be $\mathcal{I}_{\mathrm{s}}:=[0,1]$. Note that although only a subset of $\mathcal{I}_{\mathrm{inv}}$ decides to invest and thereby constitutes $\mathcal{I}_{\mathrm{s}}$, we relabel this subset to form a normalized interval $[0,1]$. We specifically examine the scenario where the sellers are homogeneous [^3] Suppliers with heterogeneous panel characteristics have been studied in, e.g., [13].. Given the aggregate capacity $c$ from (1), the homogeneity of sellers and unit length of $\mathcal{I}_{\mathrm{s}}$ imply that the panel capacity of each seller is also $c$, as $c=\bar{c}\int_{\mathcal{I}_{\mathrm{inv}}}x_{i}\,\mathrm{d}i=c\int_{\mathcal{I}_{s}}\,\mathrm{d}i$. Thus, the total solar generation from all sellers is $cG$. We now proceed with detailed modeling of the selected markets.
III-B Real-time Pricing Mechanisms
We elaborate on two alternative types of real-time pricing mechanisms proposed in the existing literature. In both mechanisms, the market price is determined by the realized values of supply and demand. The primary distinction between these two mechanisms lies in whether solar generation is treated as the same product as utility-supplied electricity. This aspect becomes crucial when consumers place additional value on renewable generation. We denote the mechanism where solar is traded as a distinct product as the product-differentiated real-time market (prt), while the other is referred to as the single-product real-time market (srt).
Within the product-differentiated real-time market, the traded product is solar electricity. Each seller $i\in\mathcal{I}_{\mathrm{s}}$ can sell any quantity of solar generation within the set $\mathcal{Q}_{i}^{\mathrm{s}}=[0,cG]$, which depends on the realized value of $G$. The seller payoff function is
$\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)=\pi q^{\mathrm{s}}_{i},$
where market equilibrium price $\pi$ and cleared quantity $q^{\mathrm{s}}_{i}$ will be defined momentarily. Given the inelastic load $L$, a buyer $i\in\mathcal{I}_{\mathrm{b}}$ can purchase any quantity of solar within the range $\mathcal{Q}^{\mathrm{b}}_{i}=[0,L]$. Any remaining demand not met by solar will be fulfilled by the utility backstop at the price $\pi_{\mathrm{u}}$. The buyer payoff function is
$\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)=(v_{i}-\pi)q^{\mathrm{b}}_{i}-\pi_{\mathrm{u}}(L-q^{\mathrm{b}}_{i}),$
where $q_{i}^{\mathrm{b}}$ is the cleared quantity for buyer $i$. Denote $q^{\mathrm{s}}:=\{q^{\mathrm{s}}_{i}\}_{i\in\mathcal{I}_{\mathrm{s}}}$ and $q^{\mathrm{b}}:=\{q^{\mathrm{b}}_{i}\}_{i\in\mathcal{I}_{\mathrm{b}}}$.
We use the following equilibrium notion to characterize the market outcome:
Definition 2 (Competitive equilibrium)
Price-allocation tuple $(\pi,q^{\mathrm{s}},q^{\mathrm{b}})$ constitutes a competitive equilibrium (CE) of the market if the following conditions are met [23]:
  • Individual rationality: Given price $\pi$, for every seller $i\in\mathcal{I}_{\mathrm{s}}$, $q_{i}^{\mathrm{s}}\in\mathop{\mathrm{argmax}}_{q\in\mathcal{Q}^{\mathrm{s}}_{i}}\Pi_{i}^{\mathrm{s}}(q,\pi)$, and for every buyer $i\in\mathcal{I}_{\mathrm{b}}$, $q_{i}^{\mathrm{b}}\in\mathop{\mathrm{argmax}}_{q\in\mathcal{Q}^{\mathrm{b}}_{i}}\Pi_{i}^{\mathrm{b}}(q,\pi)$.
  • Market clearing condition:
    $\int_{\mathcal{I}_{\mathrm{s}}}q^{\mathrm{s}}_{i}\,\,\mathrm{d}i=\int_{\mathcal{I}_{\mathrm{b}}}q^{\mathrm{b}}_{i}\,\,\mathrm{d}i.$
The CE conditions are conceptually straightforward: the resulting market outcome, comprising the equilibrium price and traded quantities, must be stable. In other words, no buyer or seller should have an incentive to unilaterally adjust their quantity, given that individual actions do not influence the market price. This price-taking assumption is consistent with our non-atomic model, where each agent is infinitesimal and cannot affect the price unilaterally.
Remark 2 (Price discovery process)
We intentionally omit the price discovery process and focus on the CE, as our interest lies in the outcomes rather than the mechanism used to obtain them. Such outcomes can arise, for instance, from buyers and sellers submitting bids to a market operator who clears the market. In our non-atomic game model, where all agents are infinitesimally small and act as price takers, it can be shown that an NE of such a bidding process corresponds to a CE that we consider here. Detailed models of price formation have been extensively studied in prior work (see references in Section I) and are beyond the scope of this paper.
With the equilibrium price $\pi$ and the cleared quantities $q^{\mathrm{s}}_{i}$ for sellers in the product-differentiated real-time market, the total expected revenue for sellers with an aggregate panel capacity $c$ over the panels’ lifespan can be calculated as
$\Pi^{\mathrm{s}}_{\mathrm{prt}}(c)=\widetilde{T}\,\mathbb{E}\left[\int_{\mathcal{I}_{\mathrm{s}}}\Pi^{\mathrm{s}}_{i}(q_{i}^{\mathrm{s}},\pi)\,\,\mathrm{d}i\right].$ (4)
We next describe the single-product real-time market, which also trades electricity. Unlike the product-differentiated real-time market, in this market both solar energy and utility-supplied energy are pooled together as a single product, regardless of the source of generation. Thus, consumers cannot express a preference or pay a premium for solar, even if they assign it a higher value. In fact, it can be shown that the outcome of this market coincides with that of the product-differentiated real-time market in the special case where $v_{i}\equiv 0$. We denote the total expected revenue for sellers in the single-product real-time market as $\Pi^{\mathrm{s}}_{\mathrm{srt}}(c)$.
III-C Contract-based Market
Unlike the real-time solar markets, the contract-based market (cb) operates well in advance of the actual delivery period. Consequently, service contracts are traded based on the expected, rather than realized, solar generation. For concreteness, we follow the terminology in [13] and focus on a setting where the contract is tied to panel capacity, meaning a panel owner rents out units of capacity and the buyer receives the realized solar output associated with that rented capacity over the duration of the contract. Since both the traded quantity and price are determined prior to the realization of generation, this market differs fundamentally from real-time pricing mechanisms. Nevertheless, for notational consistency, we will reuse the notation from the previous subsection to describe outcomes in the contract-based setting.
Since seller $i\in\mathcal{I}_{\mathrm{s}}$ can rent out any portion of their panel capacity $c$, the set of feasible amount s to sell is $\mathcal{Q}_{i}^{\mathrm{s}}=[0,c]$. Meanwhile, as buyer $i$ can in principle lease any panel capacity, we have $\mathcal{Q}_{i}^{\mathrm{b}}=\mathbb{R}_{+}$. Given the market equilibrium price $\pi$ for panel capacities, the payoff for seller $i\in\mathcal{I}_{\mathrm{s}}$ is $\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)=\pi q^{\mathrm{s}}_{i}$, and the payoff for buyer $i\in\mathcal{I}_{\mathrm{b}}$ is
$\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)=v_{i}\mathbb{E}\min\{q^{\mathrm{b}}_{i}G,L\}-\!\pi q^{\mathrm{b}}_{i}-\!\pi_{\mathrm{u}}\mathbb{E}(L\!-\!q^{\mathrm{b}}_{i}G)_{+},$ (5)
where $(z)_{+}:=\max(z,0)$, and $q_{i}^{\mathrm{s}}$ and $q_{i}^{\mathrm{b}}$ are the cleared quantities (i.e., panel capacities) for the seller and buyer, respectively.
Adapting the same CE concept as in Definition 2, with the traded product and payoff functions updated, and given the equilibrium price $\pi$ and seller cleared quantity $q^{\mathrm{s}}$, we can express the total sellers revenue for the contract-based market as
$\Pi^{\mathrm{s}}_{\mathrm{cb}}(c)=\widetilde{T}\int_{\mathcal{I}_{\mathrm{s}}}\Pi^{\mathrm{s}}_{i}(q_{i}^{\mathrm{s}},\pi)\,\,\mathrm{d}i.$ (6)
Remark 3 (Unified modeling framework)
The proposed market model serves as a general framework capable of representing a variety of solar market mechanisms, provided that the traded product and the payoff structures for buyers and sellers are appropriately defined. For example, both real-time pricing and contract-based arrangements can be formulated within this framework, with key distinctions captured through the specification of the traded product, the feasible sets of buyers and sellers, and the corresponding payoff functions. We abstract away additional implementation-specific features, as our primary objective is to investigate how different market structures shape long-term investment outcomes. While these mechanisms may differ in operational frequency, often leading to variations in price volatility as discussed in prior work [11, 12, 13], such factors do not directly affect the investment equilibrium when only expected returns matter.
Now that we have completed the modeling of these market mechanisms, we proceed to analyze how the corresponding CE can be derived for each.
IV Competitive Equilibria of Solar Markets
In this section, we derive the expected revenue for panel owners as a function of panel capacity $c$.
Lemma 1 (CE of real-time market mechanisms)
A CE for product-differentiated real-time market (prt) and single-product real-time market (srt), are listed in Table I, where $\overline{F}_{V}(\cdot):=1-F_{V}(\cdot)$ is the complementary cumulative distribution function of $v_{i}$, and $\overline{\pi}^{\mathrm{b}}_{i}:=\pi_{\mathrm{u}}+v_{i}$. Furthermore, the expected total revenue for solar panel owners under these two market mechanisms is
$\Pi^{\mathrm{s}}_{\mathrm{prt}}(c)=\widetilde{T}\,\mathbb{E}\left[\left(\pi_{\mathrm{u}}+\overline{F}^{-1}_{V}\left(\frac{cG}{L}\right)\right)\,cG\,\mathbb{1}\bigl{\{}cG\leq L\bigl{\}}\right],$ (7)
$\Pi^{\mathrm{s}}_{\mathrm{srt}}(c)=\widetilde{T}\,\mathbb{E}\left[\pi_{\mathrm{u}}\,cG\,\mathbb{1}\bigl{\{}cG\leq L\bigl{\}}\right].$ (8)
TABLE I: CE outcomes for real-time market mechanisms
Market $\pi$ $q_{i}^{\mathrm{s}}$ $q_{i}^{\mathrm{b}}$
Abundant supply
$cG>L$
prt 0 $L$ $L$
srt 0 $L$ $L$
Limited supply
$cG\leq L$
prt $\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)$ cGcG $L\mathbb{1}\{\overline{\pi}^{\mathrm{b}}_{i}\geq\pi\}$
srt $\pi_{\mathrm{u}}$ cGcG cGcG
When supply is abundant, i.e., the total realized solar generation exceeds demand, the resulting prices and allocations are identical under both types of real-time markets. This result follows from the fact that solar supply fully covers demand, rendering utility support unnecessary. As a consequence, the zero marginal cost of solar generation results in a zero equilibrium price. When the solar supply is insufficient to cover the load, the market clearing outcomes of the two market types become different. Since the utility backstop is required in this case, the market price for the single-product real-time market will be $\pi_{\mathrm{u}}$. When solar is traded as a separate product, however, we see that the scarce solar production $cG$ is allocated among the buyers with higher solar premiums first. The solar market price turns out to have a premium over the utility-supplied energy, driven by the competition among the buyers.
We now proceed to characterize the CE for the contract-based market. For each seller $i\in\mathcal{I}_{\mathrm{s}}$, the utility-maximizing choice within the feasible set for $q_{i}^{\mathrm{s}}$ is $q_{i}^{\mathrm{s}}=c$. On the buyer side, for a given market price $\pi$, let $d_{i}^{\star}(\pi)$ denote the capacity that maximizes $\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)$, as defined in (5).
To analytically characterize $d_{i}^{\star}(\pi)$, we define the expectation of a truncated version of $G$ as the following differentiable, non-increasing function:
$\mu_{G_{\mathrm{tr}}}(d_{i}):=\mathbb{E}\left[G\mathbb{1}\bigl{\{}d_{i}G\leq L\bigl{\}}\right].$ (9)
Let an extension of the inverse of this function be
[…]
where ${\mu}_{G_{\mathrm{tr}}}^{-1}(z):=\mathrm{sup}\{d_{i}|\mu_{G_{\mathrm{tr}}}(d_{i})=z\}$. We can now characterize $d_{i}^{\star}(\pi)$ as the following lemma.
Lemma 2 (Individual and aggregate demand functions under contract-based market)
The individual demand function of a buyer $i\in\mathcal{I}_{\mathrm{b}}$ is given by:
$d_{i}^{\star}(\pi):=\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}\left(\frac{\pi}{\pi_{\mathrm{u}}+v_{i}}\right).$ (10)
Accordingly, the aggregate market demand function is
$\widehat{d}^{\star}(\pi):=\int d_{i}^{\star}(\pi)\,\,\mathrm{d}i=\int\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}\left(\frac{\pi}{\pi_{\mathrm{u}}+v_{i}}\right)\,\,\mathrm{d}F_{V}(v_{i}).$ (11)
This lemma allows us to obtain the CE for the contract-based market participation game.
Lemma 3 (CE of contract-based market)
The following is a CE for the contract-based market:
$q_{i}^{\mathrm{s}}=c,\quad i\in\mathcal{I}_{\mathrm{s}};\quad q_{i}^{\mathrm{b}}=d_{i}^{\star}(\pi),\quad i\in\mathcal{I}_{\mathrm{b}},$
and $\pi$ is the solution to $c=\widehat{d}^{\star}(\pi).$ Moreover, the total revenue for solar panel owners under the contract-based market is
$\Pi^{\mathrm{s}}_{\mathrm{cb}}(c)=\widetilde{T}c\pi.$ (12)
V Equilibrium Panel Capacities
Building on the short-term market outcomes derived in the previous section, we next study the investment model to characterize the resulting equilibrium panel investment capacities.
Lemma 4 (NE panel capacities)
An NE aggregate panel capacity $c_{m}^{\mathrm{ne}}$ for each market $m\in\{\mathrm{prt},\mathrm{srt},\mathrm{cb}\}$ is a solution to the following equation:
$\Pi_{m}^{\mathrm{s}}(c)=\pi_{0}c,$ (13)
which can be specialized to individual markets as follows:
  1. 1.
    For the product-differentiated real-time market, $c_{\mathrm{prt}}^{\mathrm{ne}}$ is a solution to
    $\widetilde{T}\,\mathbb{E}\left[\left(\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right)\,G\,\mathbb{1}\bigl{\{}cG\leq L\bigl{\}}\right]=\pi_{0}.$ (14)
  2. 2.
    For the single-product real-time market, $c_{\mathrm{srt}}^{\mathrm{ne}}$ is a solution to
    $\widetilde{T}\,\mathbb{E}\left[\pi_{\mathrm{u}}\,G\,\mathbb{1}\bigl{\{}cG\leq L\bigl{\}}\right]=\pi_{0}.$ (15)
  3. 3.
    For the contract-based market,
[…]
The underlying idea behind (13) is straightforward: regardless of the specific market mechanism, the aggregate capacity reaches equilibrium when the last infinitesimal investor who decides to invest earns zero net profit, meaning $\Pi^{\mathrm{inv}}_{m,i}(x_{i},c)$ defined in (3) is zero for that investor. In other words, the equilibrium capacity can be found at the point where the capital cost curve intersects the expected revenue curve, both expressed as functions of capacity $c$. This relationship is illustrated in Fig. 2.
Figure 2: Illustration of the equilibrium condition for the panel investment problem. The relative position and shape of the curves depend on the problem parameters.
The existence of a solution to (15), and consequently to (14), is ensured under assumption (2). If the capital cost $\pi_{0}$ is too high, the expected return may fall short of covering the investment costs, even when $c$ is small. When assumption (2) is violated, the stated equilibrium condition no longer applies, and $c_{\mathrm{srt}}^{\mathrm{ne}}=0$.
While Lemma 4 enables us to use numerical simulations to compare the equilibrium capacities for each market mechanism, deeper insight comes from a theoretical comparison of the resulting aggregate capacities. To enable a more meaningful comparison, we first define a benchmark capacity as follows.
V-A Benchmark: Social Welfare Optimal Panel Investment
In addition to the market-driven investment outcomes, we also consider the social welfare optimal panel investment capacity as a benchmark, defined as the aggregate capacity at which the total payoff of consumers and investors is maximized.
To this end, we first need to define how to optimally allocate solar energy, when it is scarce, among the collection of consumers in an operation period. Given the realization of solar generation $G$, we can characterize an optimal allocation as a solution to the following optimization problem:
$\!\!\max_{\sigma:\mathcal{\mathbb{R}}_{+}\mapsto\{0,1\}}\quad$ $\int_{0}^{\bar{v}}v_{i}\sigma(v_{i})\,\,\mathrm{d}F_{V}(v_{i})$ (17a)
s.t. $\int_{0}^{\bar{v}}\sigma(v_{i})\,\,\mathrm{d}F_{V}(v_{i})\leq\min\{cG/L,1\},$ (17b)
where, for each $v_{i}$, the binary variable $\sigma(v_{i})$ indicates whether the consumer’s load is fully served by solar ($\sigma(v_{i})=1$) or not ($\sigma(v_{i})=0$). The optimal value $v^{\star}(c,G)$ then characterizes the maximum average solar premium that can be achieved given the solar capacity $c$ and the realized $G$. The social welfare optimal panel capacity is then a solution to the following optimization problem
$\max_{c\in\mathbb{R}_{+}}\enspace$ $\widetilde{T}\,\mathbb{E}\left[v^{\star}(c,G)L-\pi_{\mathrm{u}}(L-cG)_{+}\right]-\pi_{0}c.$ (18)
The following result provides an analytical characterization of this capacity.
Proposition 1 (Social welfare optimal capacity)
The social welfare optimal panel capacity, $c_{\mathrm{opt}}$, is a solution to
$\widetilde{T}\,\mathbb{E}\left[\left(\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right)\,G\,\mathbb{1}\bigl{\{}cG\leq L\bigl{\}}\right]=\pi_{0}.$ (19)
V-B Analytical Comparison of Equilibrium Capacities
As (14), (15), and (19) have similar structures, comparison among $c^{\mathrm{ne}}_{\mathrm{srt}}$, $c^{\mathrm{ne}}_{\mathrm{prt}}$, and $c_{\mathrm{opt}}$ is not difficult. However, comparing $c^{\mathrm{ne}}_{\mathrm{cb}}$ with other equilibrium capacities turns out to be challenging in general. For this purpose, we re-parametrize the solar premiums of consumers and introduce a sequence of scaled versions of the problem as follows. For each $i\in\mathcal{I}_{\mathrm{b}}$, fix $\tilde{v}_{i}$ and let $v_{i}=\epsilon\tilde{v}_{i}$ for $\epsilon\geq 0$. Given a fixed distribution $F_{\widetilde{V}}$ for $\tilde{v}_{i}$, changing the value of $\epsilon$ corresponds to shifting the distribution of the solar premiums that the population of buyers has for solar energy. While a large $\epsilon$ reflects that buyers are highly environmentally conscious and willing to pay substantial premiums for solar energy over utility electricity, a small $\epsilon$ indicates a limited willingness to do so. These scaled versions of the problem lead to the main results of this section:
Theorem 1 (Comparison of NE capacities)
The following relations hold for the equilibrium aggregate panel investment capacities induced by diverse market mechanisms.
  1. 1.
    General case: For any $\epsilon\geq 0$,
    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}.$ (20)
  2. 2.
    Small solar premium: If $\epsilon$ is in a neighborhood of 0 and $c^{\mathrm{ne}}_{\mathrm{srt}}$ is large such that
    $(1-\delta)r_{0}\leq f_{G}(g)\leq(1+\delta)r_{0},$ (21)
    holds for all $g\leq L/c^{\mathrm{ne}}_{\mathrm{srt}}$ with some $r_{0}\in\mathbb{R}_{+}$ and $\delta>0$ sufficiently small, then
    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}\lesssim c^{\mathrm{ne}}_{\mathrm{cb}},$ (22)
    where the last inequality holds in the sense that
    $c^{\mathrm{ne}}_{\mathrm{cb}}-c^{\mathrm{ne}}_{\mathrm{prt}}\geq{\color[rgb]{0,0,0}\beta\epsilon}-O\left(\epsilon^{2}\right),$ (23)
    with $\beta$ being a positive constant.
  3. 3.
    No solar premium: If $\epsilon=0$,
    $c^{\mathrm{ne}}_{\mathrm{srt}}=c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}=c^{\mathrm{ne}}_{\mathrm{cb}}.$ (24)
The first interesting result from (20) in Theorem 1 is that the single-product real-time market, in general, leads to an under-investment in solar panels relative to social welfare optimal panel capacity. In fact, with $\epsilon>0$, it is easy to identify a distribution $F_{\widetilde{V}}$ such that $c^{\mathrm{ne}}_{\mathrm{srt}}<c_{\mathrm{opt}}$, and therefore the single-product real-time energy market leads to under-investment. While one could argue that the single-product real-time market is easier to implement, particularly given the limitations of current metering infrastructure, it results in a loss of social welfare when viewed through the lens of long-term investment equilibrium.
The second observation form (20) is that the product-differentiated real-time market is guaranteed to achieve the socially optimal panel investment. This stems from the fact that the product-differentiated market indeed allocates solar energy in the same way as the social welfare optimal allocation in (17). This leads to the same expected revenue for solar panel owners and therefore the same condition for the resulting aggregate panel capacities.
When the electricity users are unwilling to pay much more for solar energy (i.e., when $\epsilon$ is small), we establish a qualitative relation between the contract-based market equilibrium capacity, $c^{\mathrm{ne}}_{\mathrm{cb}}$, and the social welfare optimal capacity, $c_{\mathrm{opt}}$. With a typical value of $\pi_{0}$ so that $c^{\mathrm{ne}}_{\mathrm{srt}}$ is relatively large, (22) suggests that the contract-based market will consistently lead to over-investment. Even though this analytical result is established only for small $\epsilon$, our numerical results suggest that the over-investment effect extends to large values of $\epsilon$ (Section VII-B). Taking a static point of view, this may be viewed as sub-optimal for the problem setting considered in this paper. However, if we incorporate potential economies of scale in panel manufacturing, over-investment may result in a smaller $\pi_{0}$ in the future and thus bring benefits to the society. Therefore, under a dynamic investment model, a policymaker may favor the contract-based market.
Finally, and perhaps most surprisingly, (24) indicates that all market mechanisms yield the same equilibrium panel capacity when users place no additional value on solar energy relative to utility-supplied electricity. This scenario may arise if utility-supplied energy is also fully renewable, or if users exhibit no specific preference for solar generation. However, given current levels of renewable penetration, non-zero values of $v_{i}$ are expected for a considerable portion of users [24]. As a result, the under/over-investment effects that market mechanisms have on the long-term solar equilibrium capacities are likely to remain substantial in the near future.
VI Heterogeneous Operation Periods
So far, we have assumed that the values/distributions of key parameters, such as loads and solar irradiation, are identical across operation periods. In practice, however, these values/distributions may vary over time due to factors such as seasonality, time of day, and day of the week. In this section, we extend our analysis to incorporate this temporal heterogeneity.
In our extended model, each period $t$ is characterized by a distinct set of parameters: the total consumer load $L_{t}$, the solar generation per unit panel $G_{t}$ with associated probability density function $f_{G_{t}}$, and the utility backstop price $\pi_{\mathrm{u}_{t}}$. Under this setup, real-time markets are cleared independently for each $t\in\mathcal{T}$, and the expected revenues of solar panel investors are determined accordingly. In contrast, contract-based markets involve long-term agreements, where consumers rent panels from suppliers once for the entire $\mathcal{T}$.
Unlike the homogeneous case, where the expected revenue is identical across periods, here the revenues vary across operation periods. Hence, we compute them individually and then take their average. As in the homogeneous case, we then scale this average by $\widetilde{T}$ to obtain the expected total revenue $\Pi_{m}^{\mathrm{s}}(c)$.
We start by characterizing the CE of short-term markets and associated revenue for investors under this new setup.
VI-1 CE of solar markets
The results from Lemma 1 and Table I, with parameters $G$, $\pi_{\mathrm{u}}$, and $L$ indexed by $t$, apply to the real-time markets in each operation period $t\in\mathcal{T}$. As a result, the expected total revenue for investors is
$\Pi^{\mathrm{s}}_{\mathrm{prt}}=\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\left[\left(\pi_{\mathrm{u}_{t}}+\overline{F}_{V}^{-1}\left(\frac{cG_{t}}{L_{t}}\right)\right)\,G_{t}\,\mathbb{1}\bigl{\{}cG_{t}\leq L_{t}\bigl{\}}\right],$ (25)
$\Pi^{\mathrm{s}}_{\mathrm{srt}}=\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\left[\pi_{\mathrm{u}_{t}}\,G_{t}\,\mathbb{1}\bigl{\{}cG_{t}\leq L_{t}\bigl{\}}\right].$ (26)
In the contract-based market, $d_{i}^{\star}(\pi)$ is a solution to
$\max_{d_{i}\geq 0}\enspace\sum_{t\in\mathcal{T}}\left[v_{i}\mathbb{E}\min\{d_{i}G_{t},L_{t}\}-\pi_{\mathrm{u}_{t}}\mathbb{E}(L_{t}-d_{i}G_{t})_{+}\right]-\pi d_{i}.$
Similar to (9), for each individual $t\in\mathcal{T}$, we can define $\mu_{G_{t,\mathrm{tr}}}(d_{i}):=\mathbb{E}\left[G_{t}\mathbb{1}\bigl{\{}d_{i}G_{t}\leq L_{t}\bigl{\}}\right]$. Using this we can analytically characterize $d_{i}^{\star}(\pi)$ as a solution to $\sum_{t\in\mathcal{T}}\left(\pi_{\mathrm{u}_{t}}+v_{i}\right)\mu_{G_{t,\mathrm{tr}}}\left(d_{i}\right)=\pi.$ The left-hand side of this equation represents the consumer’s expected total valuation for one unit of rented panel capacity over the entire planning period. We treat this quantity as a monotonically decreasing function, denoted by $w_{i}(d_{i})$, and define an extension of its inverse as
[…]
where $w_{i}^{-1}(z):=\mathrm{sup}\{d_{i}|w_{i}(d_{i})=z\}$. Thus, we get $d_{i}^{\star}(\pi)=\widetilde{w}_{i}^{-1}\left(\pi\right)$, which results in
$\widehat{d}^{\star}(\pi)=\int d_{i}^{\star}(\pi)\,\,\mathrm{d}i=\int\widetilde{w}_{i}^{-1}\left(\pi\right)\,\,\mathrm{d}i.$ (27)
Given this result, the conditions for the CE of the contract-based market given in Lemma 3 apply to this new setup. Thus, the total investors revenue is $\Pi^{\mathrm{s}}_{\mathrm{cb}}=\frac{\widetilde{T}}{T}c\pi$.
VI-2 NE panel capacities
[…]
VI-3 Social welfare optimal investment
Different operation periods may have non-identical levels of loads and expected solar generation per unit panel capacity. Thus, for each $t\in\mathcal{T}$, the optimal allocation of solar energy among consumers, denoted by $\sigma_{t}:\mathbb{R}_{+}\mapsto\{0,1\}$, and the corresponding maximum average solar premium, denoted by $v_{t}^{\star}(c,G_{t})$, should be obtained individually by solving (17) with an added subscript $t$ to parameters $L$, $G$, and function $\sigma$. Once $v_{t}^{\star}(c,G_{t})$ is obtained for all operation periods, the social welfare optimal capacity can be characterized by:
Proposition 2 (Social welfare optimal capacity with heterogeneous periods)
The social welfare optimal capacity, $c_{\mathrm{opt}}$, is a solution to
$\max_{c\in\mathbb{R}_{+}}\enspace$ $\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\!\left[v_{t}^{\star}(c,G_{t})L_{t}\!-\!\pi_{\mathrm{u}_{t}}(L_{t}-cG_{t})_{+}\right]\!-\!\pi_{0}c.$ (28)
An analytical characterization of $c_{\mathrm{opt}}$ is then obtained by solving
$\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\left[\left(\pi_{\mathrm{u}_{t}}+\overline{F}_{V}^{-1}\left(\frac{cG_{t}}{L_{t}}\right)\right)\,G_{t}\,\mathbb{1}\bigl{\{}cG_{t}\leq L_{t}\bigl{\}}\right]=\pi_{0}.$ (29)
VI-4 Analytical comparison of equilibrium capacities
Building on these results, we now extend the main findings of Theorem 1 to this new setting:
Theorem 2 (Comparison of NE capacities with heterogeneous periods)
The results in (20) and (24), corresponding to the general case and the no solar premium case of Theorem 1, remain valid in this setting. Moreover, if there exists a set of constants $\{r_{t}\}_{t\in\mathcal{T}}\in\mathbb{R}_{+}^{T}$ and a sufficiently small $\delta>0$ such that, for all $t\in\mathcal{T}$ and all $z_{t}\leq L_{t}/c^{\mathrm{ne}}_{\mathrm{srt}}$, the following condition holds
$(1-\delta)r_{t}\leq f_{G_{t}}(z_{t})\leq(1+\delta)r_{t},$ (30)
then the results associated with (22) and (23), corresponding to the small solar premium case of Theorem 1, also hold under this setting.
VII Numerical Experiments
VII-A Case Study and Problem Description
We conduct numerical experiments in this section to assess our theoretical results. Given California’s leading role in solar adoption, we focus our analysis on this state and consider a large group of potential investors in late 2023 deciding whether to install solar panels across various locations within it. To simplify the analysis, we assume a single pooled distribution-level market for the entire state and consider the three representative market frameworks discussed earlier.
We consider a 25-year planning horizon, with hourly operation periods. While our analytical results are applicable to any number of heterogeneous operation periods, we simplify the case study by grouping operation periods into two representative categories. The first group corresponds to daytime hours, between sunrise and sunset, when solar irradiation is available. The second group includes the remaining hours, during which solar irradiation is minimal or absent. This setup is equivalent to considering $\mathcal{T}=\{1,2\}$, where $t=1$ corresponds to the first group and $t=2$ to the second, with appropriate weights applied to each group in the expected revenue calculations.
Solar irradiation data from 2000 to 2022 for San Francisco, California, are used to model the distributions $f_{G_{t}}(g)$ for all $t\in\mathcal{T}$ [25]. To account for panels’ efficiency, all irradiation values are scaled by a factor of 0.2 to obtain the effective solar irradiation [26]. A kernel density estimator is then used to fit a probability density function to the gathered data, modeling the daytime irradiation density $f_{G_{1}}(g)$. The resulting density function is evaluated against the empirical distribution obtained from the histogram of the data in Fig. 3. Hours with zero or negligible effective irradiation (below $0.1~\mathrm{W}/\mathrm{m}^{2}$) are grouped into the second operation period, represented by a random variable $G_{2}$ that equals zero with probability 1.
We use survey data from [27] on consumers’ willingness to pay a premium for renewable electricity and convert these values into the per-kWh solar premium $v_{i}$ used in our model. As the data is relative to the average electricity bill for the users, we assume an average monthly household energy consumption of $600~\mathrm{kWh}$ and convert the data into solar premium for $1\,\mathrm{kWh}$ of solar energy. We also adjust the data by applying the compound inflation rate in the U.S. from 1999 to 2023. We then fit a truncated exponential density function to the data to model $f_{V}(v_{i})$, which resulted in a mean of 2.86 ¢\$/kWh and an upper bound of 16.57 ¢\$/kWh. Both $\pi_{\mathrm{u}_{1}}$ and $\pi_{\mathrm{u}_{2}}$ are set at 29 ¢\$/kWh, the average retail electricity price in California in 2023 [28]. Moreover, we set $L_{1}=27\,\mathrm{GWh}$ and $L_{2}=29\,\mathrm{GWh}$ [29]. We consider an average capital and installation cost of \$2.7 per watt [30] of solar panels, which includes labor and equipment costs.
Figure 3: Probability density function for solar irradiation
VII-B Numerical Results
VII-B1 Validity of Theorem 2
Given this setup, we first compute the expected revenues of panel investors under the real-time markets using (25) and (26). Meanwhile, for the contract-based market, we obtain the aggregate demand function, i.e., $\widehat{d}^{\star}(\pi)$, using (27). Leveraging the characterizing equations for the heterogeneous case, we then obtain the NE aggregate investment capacity under the three considered short-term markets as summarized in Table II.
TABLE II: NE aggregate investment capacity (GW)
Solar premiums scaling parameter $c_{\mathrm{srt}}^{\mathrm{ne}}$ $_{\mathrm{prt}}^{\mathrm{ne}} cc$$_{\mathrm{cb}}^{\mathrm{ne}}$ $c_{\mathrm{opt}}$
$\epsilon=1$ 67.88 70.42 71.57 70.42
$\epsilon=0$ 67.88 67.88 67.88 67.88
First, observe that this result mirrors the comparison outcome established in (20) for case 1 (general case) in Theorem 1, and likewise in Theorem 2. Moreover, although in this setup we do not scale the distribution of consumers’ premiums, i.e., $\epsilon=1$ is not in the neighborhood of zero, and assumption (30) does not hold, we still obtain the same comparison result as stated in (22), corresponding to case 2 (small solar premium) of Theorem 1, and likewise in Theorem 2. Specifically, the relation $c_{\mathrm{opt}}\lesssim c^{\mathrm{ne}}_{\mathrm{cb}}$ in that case holds here with a strict inequality, i.e., $c_{\mathrm{opt}}<c^{\mathrm{ne}}_{\mathrm{cb}}$. We then consider the scenario where consumers assign no premium to solar energy, i.e., $\epsilon=0$, which yields identical aggregate capacities across all market mechanisms, as shown in Table II and implied by (24) under case 3 (no solar premium) of Theorems 1 and 2.
In the following two subsections, we analyze how variations in consumers’ solar premiums and in panel capital and installation costs affect the comparison of equilibrium aggregate capacities across market mechanisms.
VII-B2 Sensitivity analysis with respect to consumers’ solar premiums
Beyond the specific cases of $\epsilon=0$ and $\epsilon=1$, we vary the scaling factor applied to consumers’ solar premiums and analyze the resulting investment outcomes. The results are shown in Fig. 4.
As consumers’ solar premiums increase, this figure reveals two key insights. First, the single-product real-time energy market results in a larger gap compared to the social welfare optimal capacity. This is primarily due to this market’s inability to capture the additional revenue potential from environmentally conscious consumers who are willing to pay a premium for clean energy. Second, the contract-based market consistently results in over-investment, reflected in the widening gap between $c^{\mathrm{ne}}_{\mathrm{prt}}$ and $c^{\mathrm{ne}}_{\mathrm{cb}}$.
Figure 4: Variation of equilibrium solar panel investments in response to consumers’ solar premium scaling
VII-B3 Sensitivity analysis with respect to capital and installation costs
We vary the capital and installation costs for a unit panel capacity and analyze how these changes affect the equilibrium aggregate capacities, as shown in Fig. 5.
Figure 5: Equilibrium solar panel variation with changes in installation costs
An immediate interpretation of this figure is that increasing upfront costs reduces the aggregate installed capacity of solar panels, while maintaining the inequality $c_{\mathrm{srt}}^{\mathrm{ne}}<c_{\mathrm{opt}}=c_{\mathrm{prt}}^{\mathrm{ne}}<c_{\mathrm{cb}}^{\mathrm{ne}}$ up to a certain threshold. Beyond this point, a sufficiently high $\pi_{0}$ renders solar investments unprofitable in the single-product market, causing inequality (2) to break down and resulting in $c^{\mathrm{ne}}_{\mathrm{srt}}=0$, as shown in the right-most inset subfigure of Fig 5. In contrast, due to the added revenue from environmentally conscious consumers, both $c^{\mathrm{ne}}_{\mathrm{prt}}$ and $c^{\mathrm{ne}}_{\mathrm{cb}}$ remain positive for some larger values of $\pi_{0}$. However, once $c^{\mathrm{ne}}_{\mathrm{srt}}=0$, the underlying assumptions for the theoretical results no longer hold. This breakdown is reflected in the numerical results, which indicate a shift from the expected ordering: $c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}>c^{\mathrm{ne}}_{\mathrm{cb}}>c^{\mathrm{ne}}_{\mathrm{srt}}=0$. However, given the historical decline in solar installation and capital costs and ongoing technological advances, such extreme price levels are unlikely. Thus, in most realistic cases, we expect $c_{\mathrm{cb}}^{\mathrm{ne}}>c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}>c^{\mathrm{ne}}_{\mathrm{srt}}$ to hold, consistent with our earlier theoretical insights.
VIII Concluding Remarks
This paper examines how the design of short-term electricity markets influences long-term investment in distributed solar panels, where investors’ decisions are interdependent through the expected revenues earned from participating in these markets over the panels’ lifespans. To analyze this relationship, we develop an investment model that links short-term market equilibria to the resulting long-term investment outcomes. As a key analytical tool, we introduce a unified modeling framework that derives closed-form expressions for short-term equilibrium outcomes across a broad class of market mechanisms. These expressions then serve as inputs to our investment model. Then, we focus on three representative short-term market mechanisms from the literature: a single-product real-time market, a product-differentiated real-time market, and a contract-based panel market, each reflecting a distinct approach to pricing and trading in distribution networks. For each, we apply our unified modeling framework followed by the investment model to derive analytical expressions for the corresponding long-term equilibrium panel capacity. We then theoretically establish that the product-differentiated real-time market leads to socially optimal investment, while the single-product market results in under-investment. Interestingly, when solar premiums of consumers are small, our theoretical analysis reveals that the contract-based market leads to over-investment, a pattern that our numerical results confirm even at higher levels of consumer premiums.
Future directions include integrating centralized utility planning with the distributed investment problem and incorporating policy tools, such as tax incentives or feed-in tariffs, into our analysis of long-term investment equilibria.
Appendix A Proofs for Section IV
Proof:
For each case presented in Table I, we demonstrate that the outcomes satisfy the CE conditions:
  1. 1.
    Abundant supply, prt: If $\pi=0$, any feasible $q_{i}^{\mathrm{s}}$, including $q_{i}^{\mathrm{s}}=L$, maximizes $\Pi_{i}^{\mathrm{s}}(q,0)$ for each solar owner $i\in\mathcal{I}_{\mathrm{s}}$. Moreover, $q_{i}^{\mathrm{b}}=L$ uniquely maximizes $\Pi_{i}^{\mathrm{b}}(q,0)$ for each consumer $i\in\mathcal{I}_{\mathrm{b}}$, and together these choices satisfy the market clearing condition.
  2. 2.
    Abundant supply, srt: Although $\Pi_{i}^{\mathrm{b}}(q,\pi)$ differs from the previous case, the market outcome remains the same, and the previous proof still applies.
  3. 3.
    Limited supply, prt: Given $\pi=\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)$, $q^{\mathrm{b}}_{i}$ can be obtained as
    $q^{\mathrm{b}}_{i}\in\mathop{\mathrm{argmax}}_{q\in\mathcal{Q}^{\mathrm{b}}_{i}}\left(v_{i}-\pi_{\mathrm{u}}-\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right)q-\pi_{\mathrm{u}}(L-q).$
    This is a linear optimization problem over the compact set $[0,L]$, so the optimum lies at either endpoint, determined by the slope:
    $\frac{\mathrm{d}\,\Pi_{i}^{\mathrm{b}}(q,\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right))}{\mathrm{d}q}=v_{i}-\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right).$
    If $\overline{\pi}_{i}^{\mathrm{b}}<\pi$, then $v_{i}<\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)$, so the slope is negative, and $q_{i}^{\mathrm{b}}=0$. Otherwise, the slope is positive and $q_{i}^{\mathrm{b}}=L$. At this price, the unique maximizer of $\Pi_{i}^{\mathrm{s}}(q,\pi)$ for each solar owner $i\in\mathcal{I}_{\mathrm{s}}$ is $q_{i}^{\mathrm{s}}=cG$. Given the resulting supply and demand decisions, the market clearing condition is satisfied: $\int_{\mathcal{I}_{\mathrm{s}}}cG\,\,\mathrm{d}i=\int_{\mathcal{I}_{\mathrm{b}}}L\,\mathbb{1}\left\{v_{i}\geq\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right\}\,\,\mathrm{d}i.$
  4. 4.
    Limited supply, srt: Given $\pi=\pi_{\mathrm{u}}$, the payoff for each consumer $i\in\mathcal{I}_{\mathrm{b}}$ is $\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)=-\pi_{\mathrm{u}}L$, which is independent of $q_{i}^{\mathrm{b}}$. Hence, any $q_{i}^{\mathrm{b}}\in\mathcal{Q}^{\mathrm{b}}_{i}$, including $q_{i}^{\mathrm{b}}=cG$, maximizes the consumer’s payoff. Moreover, for each solar owner $i\in\mathcal{I}_{\mathrm{s}}$, the unique maximizer of $\Pi_{i}^{\mathrm{s}}(q,\pi)$ is $q_{i}^{\mathrm{s}}=cG$. Together, these choices also satisfy the market clearing condition.
The results in Table I yield $\Pi^{\mathrm{s}}_{i}(q_{i}^{\mathrm{s}},\pi)$ for all $i\in\mathcal{I}_{\mathrm{s}}$, which can then be used in (4) to derive (7) and its special case (8).
Proof:
We begin by observing that in (5), the term $v_{i}\,\mathbb{E}\,\min\{q^{\mathrm{b}}_{i}G,L\}$ can be rewritten as $v_{i}\,L-v_{i}\,\mathbb{E}(L\!-\!q^{\mathrm{b}}_{i}G)_{+}$. Since the optimization problem in (5) is convex in $q_{i}^{\mathrm{b}}$, the first-order optimality condition is both necessary and sufficient for optimality. Taking the derivative of the objective function $\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)$ with respect to $q_{i}^{\mathrm{b}}$ and setting it to zero yields: $\pi=(\pi_{\mathrm{u}}+v_{i})\,\mu_{G_{\mathrm{tr}}}\left(q_{i}^{\mathrm{b}}\right),$ where $\mu_{G_{\mathrm{tr}}}\left(\cdot\right)$ is defined in (9). Solving for $q_{i}^{\mathrm{b}}$ yields (10), and integrating it over $i\in\mathcal{I}_{\mathrm{b}}$ gives (11). ∎
Proof:
By definition, $d_{i}^{\star}(\pi)$ belongs to the set $\arg\max_{q\in\mathcal{Q}^{\mathrm{b}}_{i}}\Pi_{i}^{\mathrm{b}}(q,\pi)$. Moreover, for any $\pi>0$, $q_{i}^{\mathrm{s}}=c$ is the unique maximizer of $\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)$, while for $\pi=0$, it is a maximizer. Given that $c=\widehat{d}^{\star}(\pi)$, these results satisfy both the individual rationality and the market clearing conditions. These results yield $\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)$ for all $i\in\mathcal{I}_{\mathrm{s}}$, which can then be used in (6) to derive (12).
Appendix B Proofs for section V
B-A Proof of Lemma 4 and Proposition 1
Proof:
Invoking (13) in (3) implies zero profit for all investors in equilibrium. Consider any investor $i\in\mathcal{I}_{\mathrm{inv}}$ with $x_{i}=0$. Since $\Pi_{m,i}^{\mathrm{inv}}(x_{i},c)=0$, unilaterally deviating to $x_{i}^{\prime}=1$ yields the same payoff by (3), and thus provides no incentive to deviate. The same holds for any investor with $x_{i}=1$. Hence, (13) captures the NE conditions of the investment game. Substituting the left-hand side of (13) with (7) and (8) yields (14) and (15), respectively. Similarly, applying (12) to the left-hand side of (13) and solving for $\pi$ gives $\pi=\pi_{0}/\widetilde{T}$. Then, using Lemma 2, which implies $c=\widehat{d}^{\star}(\pi)$, we obtain (16), completing the proof.
Proof:
First, note that $v_{i}$ for all $i\in\mathcal{I}_{\mathrm{b}}$ can be viewed as realizations of a random variable $V$ with cumulative distribution function $F_{V}$ and density $f_{V}$, as described in Section III-A. Similarly, each scaled value $\tilde{v}_{i}=\epsilon v_{i}$ corresponds to a realization of the random variable $\widetilde{V}$, with cumulative distribution function $F_{\widetilde{V}}$ and density $f_{\widetilde{V}}$.
To compute $v^{\star}(c,G)$, we solve (17), which we divide into two cases:
  1. 1.
    Abundant supply $\left(cG>L\right)$: Inequality (17b) becomes $\int_{0}^{\bar{v}}\sigma(v_{i})\,\,\mathrm{d}F_{V}(v_{i})\leq 1.$ Setting $\sigma(v_{i})=1$ for all $i\in\mathcal{I}_{\mathrm{b}}$ satisfies this constraint and maximizes the objective, since $v_{i}>0$ for all $i\in\mathcal{I}_{\mathrm{b}}$, yielding
    $v^{\star}(c,G)=\int_{0}^{\bar{v}}v_{i}\,\,\mathrm{d}F_{V}(v_{i})=\mathbb{E}[V].$ (31)
  2. 2.
    Limited supply $\left(cG\leq L\right)$: In this case, (17b) becomes $\int_{0}^{\bar{v}}\sigma(v_{i})\,\,\mathrm{d}F_{V}(v_{i})\leq cG/L$. Given $v_{i}\geq 0$ for all $i\in\mathcal{I}_{\mathrm{b}}$, a $\sigma$ that maximizes the objective is obtained by allocating the limited supply to the portion of consumers with higher solar premiums, leading to an optimal solution
    $\sigma(v_{i})=\begin{cases}1,&\mbox{if }v_{i}\geq F_{V}^{-1}\left(1-\frac{cG}{L}\right),\\ 0,&\mbox{otherwise.}\end{cases}$
    This solution yields the corresponding optimal value of (17a) as $v^{\star}(c,G)=\int_{F_{V}^{-1}\left(1-\frac{cG}{L}\right)}^{\bar{v}}v_{i}\,\,\mathrm{d}F_{V}(v_{i})$.
Substituting the expressions for $v^{\star}(c,G)$ into the optimization problem (18) yields:
$\max_{c\in\mathbb{R}_{+}}\quad\widetilde{T}\,\mathbb{E}\Big{[}\,$ $L\,\mathbb{E}[V]-L\,\mathbb{1}\{cG\leq L\}\int_{0}^{F_{V}^{-1}\left(1-\frac{cG}{L}\right)}\!v_{i}\,\,\mathrm{d}F_{V}(v_{i})$
$-\pi_{\mathrm{u}}(L-cG)_{+}\,\Big{]}-\pi_{0}c.$
Under our assumptions, we can verify that the second derivative of the objective with respect to $c$ is given by
$\int_{0}^{L/c}\frac{g^{2}}{L}\,\frac{f_{G}(g)}{f_{V}\left(F_{V}^{-1}\left(1-\tfrac{cg}{L}\right)\right)}\,\mathrm{d}g\leq 0,$
which implies that the objective function is concave in $c$. Thus, the problem is convex, and the first-order optimality condition is sufficient to characterize the global optimum. Applying this condition yields the following analytical characterization of $c_{\mathrm{opt}}$ as the solution to
$\widetilde{T}\,\mathbb{E}\left[\left(\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right)G\,\mathbb{1}\{cG\leq L\}\right]=\pi_{0},$
which completes the proof. ∎
B-B Proof of Theorem 1
We establish each case individually:
Proof:
Since (14) and (19) are identical, we have $c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}$. Furthermore, for any $\epsilon\geq 0$, $c^{\mathrm{ne}}_{\mathrm{prt}}$ solves the modified version of (14):
$\pi_{0}-\widetilde{T}\,\mathbb{E}\left[\left(\pi_{\mathrm{u}}+\epsilon\,\overline{F}_{\widetilde{V}}^{-1}\left(\frac{cG}{L}\right)\right)G\,\mathbb{1}\{cG\leq L\}\right]=0,$ (32)
which is equivalent to
$\int_{0}^{L/c}\left(\pi_{\mathrm{u}}+\epsilon\,\overline{F}_{\widetilde{V}}^{-1}\left(\frac{cg}{L}\right)\right)g\,f_{G}(g)\,\mathrm{d}g=\frac{\pi_{0}}{\widetilde{T}}$. (33)
Similarly, $c^{\mathrm{ne}}_{\mathrm{srt}}$ is characterized as a solution to an equivalent form of (15):
$\int_{0}^{L/c}\pi_{\mathrm{u}}\,g\,f_{G}(g)\,\mathrm{d}g=\frac{\pi_{0}}{\widetilde{T}}$. (34)
Since the terms $\pi_{\mathrm{u}},\epsilon\,\overline{F}_{\widetilde{V}}^{-1}(\cdot)$, and $g$ are all non-negative, we have
$\left(\pi_{\mathrm{u}}+\epsilon\,\overline{F}_{\widetilde{V}}^{-1}\left(\frac{cg}{L}\right)\right)g\,f_{G}(g)\geq\pi_{\mathrm{u}}g\,f_{G}(g).$
Therefore, by comparing (33) and (34), we conclude that $\frac{L}{c^{\mathrm{ne}}_{\mathrm{srt}}}\geq\frac{L}{c^{\mathrm{ne}}_{\mathrm{prt}}}$, which implies that $c^{\mathrm{ne}}_{\mathrm{prt}}\geq c^{\mathrm{ne}}_{\mathrm{srt}}$. ∎
Proof:
In this case, i.e., $\epsilon=0$,  (33) and (34) are identical. Therefore, $c_{\mathrm{opt}}=c^{\mathrm{ne}}_{\mathrm{prt}}=c^{\mathrm{ne}}_{\mathrm{srt}}$, and it remains to show $c^{\mathrm{ne}}_{\mathrm{cb}}=c^{\mathrm{ne}}_{\mathrm{srt}}$.
[…]
$\max_{d_{i}\geq 0}\quad\epsilon\tilde{v}_{i}\,\mathbb{E}\min\{d_{i}G,L\}-\frac{\pi_{0}}{\widetilde{T}}d_{i}-\pi_{\mathrm{u}}\,\mathbb{E}(L-d_{i}G)_{+}.$
[…]
$\left(\pi_{\mathrm{u}}+\epsilon\tilde{v}_{i}\right)\mathbb{E}\left[G\,\mathbb{1}\left\{d_{i}^{\star}(\pi)G\leq L\right\}\right]+\frac{\pi_{0}}{\widetilde{T}}=0.$ (35)
[…]
Proof:
Define $c_{0}:=c_{\mathrm{srt}}^{\mathrm{ne}}$. The main steps of the proof are summarized as follows: 1) we treat the left-hand sides of (32) and (35) as functions of both $\epsilon$ (via its influence on $v$) and $c$; 2) we show that if $\epsilon$ lies in a neighborhood of 0, then the corresponding capacities (i.e., the values of $c$ that solve (32) and (35)) must lie in a neighborhood of $c_{0}$; 3) we derive the first-order Taylor expansions of these functions at $(\epsilon,c)=(0,c_{0})$, and provide explicit bounds on the higher-order terms; and 4) we compare the resulting Taylor expansions term by term.
We defer the detailed proofs of the first three steps to Appendix C, and summarize their results below in the form of lemmas.
Lemma 5 (First-order expansion of $c_{\mathrm{prt}}^{\mathrm{ne}}$ around $\epsilon=0$)
For a sufficiently small $\epsilon>0$, we have
$c_{\mathrm{prt}}^{\mathrm{ne}}=c_{0}-\frac{\epsilon\,\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{0}G}{L}\right)\,G\,\mathbb{1}\bigl{\{}c_{0}G\leq L\bigl{\}}\right]}{\pi_{\mathrm{u}}\mu_{G_{\mathrm{tr}}}^{\prime}(c_{0})}+O\left(\epsilon^{2}\right),$ (36)
where $\mu_{G_{\mathrm{tr}}}(c_{0})$ is defined in (9).
Lemma 6 (First-order expansion of $c_{\mathrm{cb}}^{\mathrm{ne}}$ around $\epsilon=0$)
For a sufficiently small $\epsilon>0$, we have
$d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)=c_{0}-\frac{\epsilon\tilde{v}_{i}\mu_{G_{\mathrm{tr}}}(c_{0})}{\pi_{\mathrm{u}}\mu_{G_{\mathrm{tr}}}^{\prime}\left(c_{0}\right)}+O\left(\epsilon^{2}\right),$ (37)
and consequently
$c_{\mathrm{cb}}^{\mathrm{ne}}=c_{0}-\frac{\epsilon\mathbb{E}[\widetilde{V}]\mu_{G_{\mathrm{tr}}}(c_{0})}{\pi_{\mathrm{u}}\mu_{G_{\mathrm{tr}}}^{\prime}(c_{0})}+O\left(\epsilon^{2}\right).$ (38)
Equipped with these lemmas, we can compare $c_{\mathrm{cb}}^{\mathrm{ne}}$ and $c_{\mathrm{prt}}^{\mathrm{ne}}$ by isolating the non-identical terms in (36) and (38).
Note that as $\mu_{G_{\mathrm{tr}}}^{\prime}(c_{0})=\left(-L^{2}/\left(c_{0}\right)^{3}\right)\,f_{G}\left(L/c_{0}\right)<0$, the proof is complete if we show
$\underbrace{\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{0}G}{L}\right)G\,\mathbb{1}\left\{c_{0}G\leq L\right\}\right]}_{:=A}<\underbrace{\mathbb{E}[\widetilde{V}]\,\mu_{G_{\mathrm{tr}}}(c_{0})}_{:=B}.$ (39)
For notational simplicity, let $\tilde{v}(p):=\overline{F}_{\widetilde{V}}^{-1}(p)$ denote a decreasing function that maps $[0,1]$ to $0,\bar{v}]$. Thus,
$A=\mathbb{E}\left[\tilde{v}\left(\frac{c_{0}G}{L}\right)G\,\mathbb{1}\{c_{0}G\leq L\}\right].$
Since $\tilde{v}(1)=0$, we have $\tilde{v}\left(\frac{c_{0}G}{L}\right)=\int_{c_{0}G/L}^{1}-\tilde{v}^{\prime}(p)\,\mathrm{d}p,$ which implies
$A=\int_{0}^{1}{\color[rgb]{0,0,0}\left(-\tilde{v}^{\prime}(p)\right)}\,\mathbb{E}\left[G\,\mathbb{1}\bigl{\{}\frac{c_{0}G}{L}\leq p\bigl{\}}\right]\,\,\mathrm{d}p.$ (40)
Observe that
$\mathbb{E}[\widetilde{V}]=-\int_{0}^{\bar{v}}\tilde{v}_{i}\,\mathrm{d}\overline{F}_{\widetilde{V}}(\tilde{v}_{i})=\int_{0}^{\bar{v}}\overline{F}_{\widetilde{V}}(\tilde{v}_{i})\,\,\mathrm{d}\tilde{v}_{i},$ (41)
where the first equality follows from the definition of expectation and the identity $\mathrm{d}F_{\widetilde{V}}=-\mathrm{d}\overline{F}_{\widetilde{V}}$, and the second equality follows from integration by parts.
Next, recall that by the integral identity for inverse functions, if $f:[a,b]\to\mathbb{R}$ is strictly monotonic, then the following identity holds:
$\int_{a}^{b}f(x)\,dx=bf(b)-af(a)-\int_{f(a)}^{f(b)}f^{-1}(y)\,\,dy.$
Since $\overline{F}_{\widetilde{V}}$ is strictly monotonic, we can apply this identity and obtain:
$\int_{0}^{\bar{v}}\overline{F}_{\widetilde{V}}(\tilde{v}_{i})\,\mathrm{d}\tilde{v}_{i}=$ $\left(\bar{v}\right)\underset{0}{\underbrace{\overline{F}_{\widetilde{V}}(\bar{v})}}-0-\int_{\overline{F}_{\widetilde{V}}(0)}^{\overline{F}_{\widetilde{V}}(\bar{v})}\overline{F}_{\widetilde{V}}^{-1}\left(p\right)\,\,\mathrm{d}p$
$=\int_{\overline{F}_{\widetilde{V}}(\bar{v})}^{\overline{F}_{\widetilde{V}}(0)}\underset{\tilde{v}(p)}{\underbrace{\overline{F}_{\widetilde{V}}^{-1}\left(p\right)}}\,\,\mathrm{d}p=\int_{0}^{1}\tilde{v}(p)\,p\,\,\mathrm{d}p{\color[rgb]{0,0,0}.}$ (42)
By integration by parts, we have
$\int_{0}^{1}\tilde{v}(p)\,p\,\,\mathrm{d}p=-\int_{0}^{1}\tilde{v}^{\prime}(p)\,p\,\,\mathrm{d}p.$ (43)
Combining (41), (42), and (43), we obtain
$\mathbb{E}[\widetilde{V}]=-\int_{0}^{1}\tilde{v}^{\prime}(p)\,p\,\,\mathrm{d}p.$
Substituting this result into $B$ then gives
$B=\int_{0}^{1}{\color[rgb]{0,0,0}\left(-v^{\prime}(p)\right)}\,\mathbb{E}\left[p\,G\,\mathbb{1}\bigl{\{}\frac{c_{0}G}{L}\leq 1\bigl{\}}\right]\,\,\mathrm{d}p.$ (44)
It follows from (40) and (44) that, since $\tilde{v}(p)<0$, establishing $A<B$ reduces to showing:
$\mathbb{E}\left[G\,\mathbb{1}\left\{\frac{c_{0}G}{L}\leq p\right\}\right]<\mathbb{E}\left[p\,G\,\mathbb{1}\left\{\frac{c_{0}G}{L}\leq 1\right\}\right].$
Note that
$\mathbb{E}\left[G\,\mathbb{1}\left\{\frac{c_{0}G}{L}\leq p\right\}\right]=\int_{0}^{Lp/c_{0}}g\,f_{G}(g)\,\,\mathrm{d}g.$
Furthermore, under assumption (21), we have $\int_{0}^{Lp/c_{0}}gf_{G}(g)\,\,\mathrm{d}g\leq(1+\delta)r_{0}\frac{1}{2}\left(\frac{Lp}{c_{0}}\right)^{2}$, which leads to the following upper bound for AA:
$A\leq(1+\delta)r_{0}\frac{1}{2}\left(\frac{L}{c_{0}}\right)^{2}\int_{0}^{1}-\tilde{v}^{\prime}(p)\,p^{2}\,\,\mathrm{d}p.$
Similarly, $\mathbb{E}\left[p\,G\,\mathbb{1}\left\{\frac{c_{0}G}{L}\leq 1\right\}\right]=\int_{0}^{L/c_{0}}p\,g\,f_{G}(g)\,\,\mathrm{d}g$, and applying assumption (21) again, we obtain the lower bound for $B$:
$B\geq(1-\delta)r_{0}\frac{1}{2}\left(\frac{L}{c_{0}}\right)^{2}\int_{0}^{1}-\tilde{v}^{\prime}(p)\,p\,\,\mathrm{d}p.$
We can proceed by defining $\lambda:=\frac{\int_{0}^{1}-\tilde{v}^{\prime}(p)\,p^{2}\,\mathrm{d}p}{\int_{0}^{1}-\tilde{v}^{\prime}(p)\,p\,\mathrm{d}p}$, where $0<\lambda<1$ since $p\in[0,1]$. One can find a $\delta>0$ such that
$\frac{1+\delta}{1-\delta}\lambda\leq\frac{\lambda+1}{2}.$ (45)
Multiplying both sides of (45) by $\frac{r_{0}}{2}\left(\frac{L}{c_{0}}\right)^{2}$ and applying the derived bounds for AA and $B$ yields:
$A\leq\frac{\lambda+1}{2}B{\color[rgb]{0,0,0}<B},$
which proves (39). Thus, we conclude:
$\mathbb{E}[\widetilde{V}]\mu_{G_{\mathrm{tr}}}(c_{0})\geq\frac{2}{\lambda+1}\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{0}G}{L}\right)\,G\,\mathbb{1}\bigl{\{}c_{0}G\leq L\bigl{\}}\right],$
which can be used in (36) and (38) to derive (23), where
$\beta:=\frac{\left(1-\lambda\right)\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{0}G}{L}\right)\,G\,\mathbb{1}\bigl{\{}c_{0}G\leq L\bigl{\}}\right]}{-\left(1+\lambda\right)\pi_{\mathrm{u}}\mu_{G_{\mathrm{tr}}}^{\prime}(c_{0})}>0.$
Appendix C Proof of Lemma 5 and Lemma 6
C-A Proof of Lemma 5
We can use (9) to reformulate (32) as
$\underbrace{\frac{\pi_{0}}{\widetilde{T}}-\pi_{\mathrm{u}}\,\mu_{G_{\mathrm{tr}}}(c^{\mathrm{ne}}_{\mathrm{prt}})-\epsilon\,k(c^{\mathrm{ne}}_{\mathrm{prt}})}_{:=\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})}=0,$ (46)
where
$k(c^{\mathrm{ne}}_{\mathrm{prt}}):=\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{\mathrm{prt}}^{\mathrm{ne}}G}{L}\right)\,G\,\mathbb{1}\bigl{\{}c^{\mathrm{ne}}_{\mathrm{prt}}G\leq L\bigl{\}}\right].$
For given values of $\pi_{0}$, $\pi_{\mathrm{u}}$, and $\epsilon$, the solution to $\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})=0$ determines $c^{\mathrm{ne}}_{\mathrm{prt}}$. We first show that if $\epsilon$ is sufficiently close to 0, then the corresponding $c^{\mathrm{ne}}_{\mathrm{prt}}$ lies in a neighborhood of $c_{0}$.
To emphasize the dependence of the solution $c^{\mathrm{ne}}_{\mathrm{prt}}$ on $\epsilon$, we write $c^{\mathrm{ne}}_{\mathrm{prt}}=\Psi(\epsilon)$, where $\Psi:[0,1]\to\mathbb{R}$. From (24), it follows that $\Psi(0)=c_{0}$. Furthermore, the following lemma shows that $\Psi(\epsilon)$ is Lipschitz continuous with constant $L_{\Psi}$.
Lemma 7 (Lipschitz continuity of $\Psi(\epsilon)$ near $\epsilon=0$)
In a neighborhood of $\epsilon=0$, the function $\Psi(\epsilon)$ is Lipschitz continuous with Lipschitz constant $L_{\Psi}$.
Equipped with Lemma 7, we have $|\Psi(\epsilon)-\Psi(0)|\leq L_{\Psi}|\epsilon-0|,$ which implies
$|c^{\mathrm{ne}}_{\mathrm{prt}}-c_{0}|\leq L_{\Psi}\epsilon.$ (47)
In other words, if $\epsilon$ is sufficiently close to 0, then $c^{\mathrm{ne}}_{\mathrm{prt}}$ remains close to $c_{0}$. This proximity allows us to expand $\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})$ in a Taylor series at $(0,c_{0})$:
$\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})=$ $-\epsilon\,k\left(c_{0}\right)-\pi_{\mathrm{u}}\,(c^{\mathrm{ne}}_{\mathrm{prt}}-c_{0})\,\mu_{G_{\mathrm{tr}}}^{\prime}(c_{0})$
$+O(\lVert(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})-(0,c_{0})\rVert^{2})=0.$ (48)
Using (47), we can bound the higher-order terms as
$\lVert(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})-(0,c_{0})\rVert^{2}\leq\epsilon^{2}\left(1+L_{\Psi}^{2}\right),$
which can be applied in (C-A) to reformulate it as
$\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})=$ $-\epsilon\,k\left(c_{0}\right)-\pi_{\mathrm{u}}\,(c^{\mathrm{ne}}_{\mathrm{prt}}-c_{0})\,\mu_{G_{\mathrm{tr}}}^{\prime}(c_{0})$
$+O(\epsilon^{2})=0.$ (49)
Isolating $c^{\mathrm{ne}}_{\mathrm{prt}}$ then yields (36). It remains to prove Lemma 7.
Proof:
We first show that there exists a $\Delta>0$ such that for all $c_{1},c_{2}\in[c_{0}-\Delta,c_{0}+\Delta]$ with $c_{1}>c_{2}$, the following holds for some $m,M\in\mathbb{R}_{++}$:
$0<m\leq\frac{\phi_{\mathrm{prt}}(\epsilon,c_{1})-\phi_{\mathrm{prt}}(\epsilon,c_{2})}{c_{1}-c_{2}}\leq M.$ (50)
According to (46), this inequality is equivalent to:
$0<m\leq$ $\pi_{\mathrm{u}}\left(\frac{\mu_{G_{\mathrm{tr}}}(c_{2})-\mu_{G_{\mathrm{tr}}}(c_{1})}{c_{1}-c_{2}}\right)$
$+\epsilon\left(\frac{k(c_{2})-k(c_{1})}{c_{1}-c_{2}}\right)\leq M.$ (51)
To prove the inequality, observe that for all $c\in[c_{0}-\Delta,c_{0}+\Delta]$, there exist some $m_{1},M_{1}\in\mathbb{R}_{++}$ such that
$M_{1}\leq\pi_{\mathrm{u}}\mu_{G_{\mathrm{tr}}}^{\prime}(c)=\pi_{\mathrm{u}}\frac{-L^{2}}{c^{3}}f_{G}\left(\frac{L}{c}\right)\leq-m_{1}<0.$ (52)
By the Mean Value Theorem [31], there exists a point $c_{3}\in(c_{2},c_{1})$ such that:
$\frac{\mu_{G_{\mathrm{tr}}}(c_{2})-\mu_{G_{\mathrm{tr}}}(c_{1})}{c_{1}-c_{2}}=-\frac{\mu_{G_{\mathrm{tr}}}(c_{1})-\mu_{G_{\mathrm{tr}}}(c_{2})}{c_{1}-c_{2}}=-\mu_{G_{\mathrm{tr}}}^{\prime}(c_{3}).$
Therefore, for all $c_{1},c_{2}\in[c_{0}-\Delta,c_{0}+\Delta]$ with $c_{1}>c_{2}$, it follows that
$0<m_{1}\leq\pi_{\mathrm{u}}\left(\frac{\mu_{G_{\mathrm{tr}}}(c_{2})-\mu_{G_{\mathrm{tr}}}(c_{1})}{c_{1}-c_{2}}\right)\leq M_{1}.$ (53)
Next, we consider the derivative of $k(c)$, which is given by:
$k^{\prime}(c)=-\mathbb{E}\left[\frac{G^{2}}{L}\frac{\mathbb{1}\bigl{\{}cG\leq L\bigl{\}}}{f_{\widetilde{V}}\left(\overline{F}_{\widetilde{V}}^{-1}\left(\frac{cG}{L}\right)\right)}\right]<0.$
For all $c\in[c_{0}-\Delta,c_{0}+\Delta]$ and $\epsilon>0$, there exist constants $m_{2},M_{2}\in\mathbb{R}_{++}$ such that:
$M_{2}\leq\epsilon k^{\prime}(c)\leq-m_{2}<0.$ (54)
Then, by the Mean Value Theorem, it follows that:
$0<m_{2}\leq\epsilon\left(\frac{k(c_{2})-k(c_{1})}{c_{1}-c_{2}}\right)\leq M_{2}.$ (55)
Adding inequalities (53) and (55), and letting $m=m_{1}+m_{2}$ and $M=M_{1}+M_{2}$, we obtain (C-A), which in turn implies (50).
Next, observe that $\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}}):[0,1]\times\mathbb{R}\to\mathbb{R}$ is continuous, and $[0,1]$ is a compact set. These properties, together with the bounded derivative condition in (50), ensure that $\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})$ satisfies the conditions of the Lipschitz Implicit Function Theorem (see Theorem 6 of [32]) in the considered neighborhood of $c_{0}$. Thus, there exists a unique function $\Psi:[0,1]\to\mathbb{R}$ such that $\phi_{\mathrm{prt}}(\epsilon,\Psi(\epsilon))=0$, and $\Psi$ is continuous. To establish that $\Psi$ is Lipschitz continuous, it suffices to show that there exists a constant $L_{\Psi}\geq 0$ such that for all $\epsilon_{1},\epsilon_{2}\in[0,1]$,
$\frac{|\Psi(\epsilon_{2})-\Psi(\epsilon_{1})|}{|\epsilon_{2}-\epsilon_{1}|}\leq L_{\Psi}.$
Let $c_{1}=\Psi(\epsilon_{1})$ and $c_{2}=\Psi(\epsilon_{2})$. From (46), it follows that they satisfy $\frac{\pi_{0}}{\widetilde{T}}-\pi_{\mathrm{u}}\,\mu_{G_{\mathrm{tr}}}(c_{1})-\epsilon_{1}\,k(c_{1})=0$ and $\frac{\pi_{0}}{\widetilde{T}}-\pi_{\mathrm{u}}\,\mu_{G_{\mathrm{tr}}}(c_{2})-\epsilon_{2}\,k(c_{2})=0$, respectively. Subtracting the second equation from the first gives:
$\pi_{\mathrm{u}}\left[\mu_{G_{\mathrm{tr}}}(c_{2})-\mu_{G_{\mathrm{tr}}}(c_{1})\right]+\epsilon_{2}k(c_{2})-\epsilon_{1}k(c_{1})=0.$ (56)
Define the function $H:[0,1]\to\mathbb{R}$ as $H(\tau):=\,\left(\epsilon_{1}+\tau(\epsilon_{2}-\epsilon_{1})\right)k\left(c_{1}+\tau(c_{2}-c_{1})\right)+\pi_{\mathrm{u}}\,\mu_{G_{\mathrm{tr}}}\left(c_{1}+\tau(c_{2}-c_{1})\right).$ Then, (56) is equivalent to H(1)H(0)=0H(1)-H(0)=0.
By applying the Mean Value Theorem to the left-hand side of this equation, we obtain:
$H(1)-H(0)=H^{\prime}(\zeta)=0,$
for some $\zeta\in(0,1)$. Moreover, we have
$H^{\prime}(\zeta)=\,$ $\pi_{\mathrm{u}}\,\mu_{G_{\mathrm{tr}}}^{\prime}\left(c_{1}+\zeta(c_{2}-c_{1})\right)(c_{2}-c_{1})$
$+(\epsilon_{2}-\epsilon_{1})\,k\left(c_{1}+\zeta(c_{2}-c_{1})\right)$
$+\left(\epsilon_{1}+\zeta(\epsilon_{2}-\epsilon_{1})\right)\,k^{\prime}\left(c_{1}+\zeta(c_{2}-c_{1})\right)(c_{2}-c_{1}).$
Since $H^{\prime}(\zeta)=0$, it follows that
$\frac{c_{2}-c_{1}}{\epsilon_{2}-\epsilon_{1}}=\frac{k\left(c_{1}+\zeta(c_{2}-c_{1})\right)}{J},$ (57)
where
$J:=$ $-\pi_{\mathrm{u}}\,\mu_{G_{\mathrm{tr}}}^{\prime}\left(c_{1}+\zeta(c_{2}-c_{1})\right)$
$-\left(\epsilon_{1}+\zeta(\epsilon_{2}-\epsilon_{1})\right)k^{\prime}\left(c_{1}+\zeta(c_{2}-c_{1})\right).$
It remains to upper bound the absolute value of the numerator and lower bound $|J|$ on the right-hand side of (57). To this end, note that
$\left|k\left(c_{1}+\zeta(c_{2}-c_{1})\right)\right|\leq\bar{v}\,\mathbb{E}[G].$
Moreover, from (52), the first term in $J$ is bounded below by a positive constant $m_{1}$, and from (54), the second term is similarly bounded below by a positive constant $m_{2}$. Thus, we conclude:
$\frac{|c_{2}-c_{1}|}{|\epsilon_{2}-\epsilon_{1}|}=\frac{\left|k\left(c_{1}+\zeta(c_{2}-c_{1})\right)\right|}{|J|}\leq\frac{\bar{v}\,\mathbb{E}[G]}{\left(m_{1}+m_{2}\right)}=L_{\Psi},$
which completes the proof of Lemma 7.
C-B Proof of Lemma 6
Using (9), we can reconstruct (35) as:
$\underbrace{\frac{\pi_{0}}{\widetilde{T}}-\pi_{\mathrm{u}}\,\mu_{G_{\mathrm{tr}}}\left(d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)\right)-\epsilon\tilde{v}_{i}\,\mu_{G_{\mathrm{tr}}}\left(d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)\right)}_{:=\phi_{\mathrm{cb},{i}}\left(\epsilon,d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)\right)}=0.$
[…]
$\left\lvert d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)-c_{0}\right\rvert=\left\lvert\Delta_{\mathrm{cb},i}\left(\epsilon\right)\right\rvert$,
where
$\Delta_{\mathrm{cb},i}\left(\epsilon\right):=\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}\left(\frac{\pi_{0}}{\widetilde{T}\left(\pi_{\mathrm{u}}+\epsilon\tilde{v}_{i}\right)}\right)-\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}\left(\frac{\pi_{0}}{\widetilde{T}\,\pi_{\mathrm{u}}}\right),$ (58)
and it remains to bound $\lvert\Delta_{\mathrm{cb},i}\left(\epsilon\right)\rvert$, which diminishes as $\epsilon\rightarrow 0$.
To proceed, define the function $r_{i}(\epsilon):=\frac{\pi_{0}}{\widetilde{T}\left(\pi_{\mathrm{u}}+\epsilon\tilde{v}_{i}\right)}$. Then, $\Delta_{\mathrm{cb},i}\left(\epsilon\right)=\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(r_{i}(\epsilon))-\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(r_{i}(0)).$ We now show that $\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(r_{i}(\cdot))$ is Lipschitz continuous by proving that both $r_{i}(\cdot)$ and $\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(\cdot)$ are individually Lipschitz continuous. Note that both of these functions are decreasing, continuous, and differentiable on their respective domain. Thus, the Lipschitz constant for $r_{i}(\epsilon)$ is given by $L_{r_{i}}=\max_{\epsilon}\left|\frac{\mathrm{d}r_{i}\left(\epsilon\right)}{\mathrm{d}\epsilon}\right|=\frac{\pi_{0}\tilde{v}_{i}}{\widetilde{T}\,\pi_{\mathrm{u}}^{2}}.$ Similarly, the Lipschitz constant for $\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(\cdot)$ is obtained as
$L_{\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}}=\max_{y}\quad\left|\frac{\mathrm{d}\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}\left(y\right)}{\mathrm{d}y}\right|=\max_{y}\quad\left|\frac{\big{(}\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(y)\big{)}^{3}}{L^{2}f_{G}\left(\frac{L}{\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(y)}\right)}\right|,$
where $\frac{\pi_{0}}{\widetilde{T}\left(\pi_{\mathrm{u}}+\bar{v}\right)}\leq y\leq\frac{\pi_{0}}{\widetilde{T}\,\pi_{\mathrm{u}}}$, and therefore $\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(y)$ is bounded as $\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}\left(\frac{\pi_{0}}{\widetilde{T}\,\pi_{\mathrm{u}}}\right)\leq\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(y)\leq\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}\left(\frac{\pi_{0}}{\widetilde{T}\left(\pi_{\mathrm{u}}+\bar{v}\right)}\right)$. Finally, since the composition of two Lipschitz continuous functions is also Lipschitz continuous, the function $\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(r_{i}(\epsilon))$ is Lipschitz continuous with constant $L_{\mathrm{cb},i}=L_{\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}}\,L_{r_{i}}$. Therefore,
$\frac{\left|\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(r_{i}(\epsilon))-\tilde{\mu}_{G_{\mathrm{tr}}}^{-1}(r_{i}(0))\right|}{\epsilon}\leq L_{\mathrm{cb},i},$
and combining this with (58), we conclude that
$0\leq\Delta_{\mathrm{cb},i}=\left|\Delta_{\mathrm{cb},i}\right|\leq\epsilon\,L_{\mathrm{cb},i}$ (59)
where the first inequality holds because $\Delta_{\mathrm{cb},i}(r_{i}(\epsilon))$ is an increasing function with respect to $\epsilon$.
[…]
$-\epsilon\tilde{v}_{i}\mu_{G_{\mathrm{tr}}}(c_{0})$
$+O\left(\left\lVert\left(\epsilon,d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)\right)-(0,c_{0})\right\rVert^{2}\right)$
$=0.$ (60)
We have
$\left\lVert\left(\epsilon,d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)\right)-(0,c_{0})\right\rVert^{2}=\epsilon^{2}+\underbrace{\left(d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)-c_{0}\right)^{2}}_{\left(\Delta_{\mathrm{cb},i}\right)^{2}}.$
Then, based on (59), we obtain
$\left\lVert\left(\epsilon,d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)\right)-(0,c_{0})\right\rVert^{2}\leq\epsilon^{2}\left(1+\left(L_{\mathrm{cb},i}\right)^{2}\right).$
Using this result in (C-B), we can reformulate (C-B) as
[…]
$-\epsilon\tilde{v}_{i}\mu_{G_{\mathrm{tr}}}(c_{0})+O\!\left(\epsilon^{2}\right)=0.$ (61)
[…]
Appendix D Proofs for Section VI
D-A Proof of Proposition 2
The expression for $v_{t}^{\star}(c,G_{t})$ is derived individually for each $t\in\mathcal{T}$ using the same method as in the proof of Proposition 1. Substituting these expressions into (28) yields:
$\max_{c\in\mathbb{R}_{+}}\enspace\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\Biggr{[}$ $L_{t}\mathbb{1}\!\left\{cG_{t}\leq L_{t}\right\}\!\int_{0}^{F_{V}^{-1}\left(1-\frac{cG_{t}}{L_{t}}\right)}v_{i}\,\,\mathrm{d}F_{V}(v_{i})$
$+L_{t}\mathbb{E}[V]-\pi_{\mathrm{u_{t}}}\mathbb{(}L_{t}-cG_{t})_{+}\Biggr{]}-\pi_{0}c.$
The objective is concave in $c$ since its second derivative is given by
$\sum_{t\in\mathcal{T}}\int_{0}^{L_{t}/c}\frac{\left(g_{t}\right)^{2}}{L_{t}}\,\frac{f_{G}\left(g_{t}\right)}{f_{V}\left(F_{V}^{-1}\left(1-\tfrac{cg_{t}}{L_{t}}\right)\right)}\,\,\mathrm{d}g_{t}\leq 0,$
which ensures that the problem is convex. Therefore, the first-order optimality condition is sufficient to analytically characterize $c_{\mathrm{opt}}$ as (29), completing the proof.         $\blacksquare$
D-B Proof of Theorem 2
We prove each case individually:
Proof:
Since the characterizing equation for $c_{\mathrm{prt}}^{\mathrm{ne}}$ is identical to that of (29), it follows that $c_{\mathrm{prt}}^{\mathrm{ne}}=c_{\mathrm{opt}}$. Also, for any $\epsilon\geq 0$, $c_{\mathrm{prt}}^{\mathrm{ne}}$ is a solution to
$\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\left[\left(\pi_{\mathrm{u}_{t}}+\epsilon\overline{F}_{\widetilde{V}}^{-1}\left(\frac{cG_{t}}{L_{t}}\right)\right)\,G_{t}\,\mathbb{1}\bigl{\{}cG_{t}\leq L_{t}\bigl{\}}\right]=\pi_{0},$ (62)
which can be rewritten as
$\sum_{t\in\mathcal{T}}\int_{0}^{L_{t}/c}\left(\pi_{\mathrm{u}_{t}}+\epsilon\overline{F}_{\widetilde{V}}^{-1}\left(\frac{cg_{t}}{L_{t}}\right)\right)g_{t}f_{G_{t}}(g_{t})\,\,\mathrm{d}g_{t}=\pi_{0}\frac{T}{\widetilde{T}}$. (63)
Similarly, $c^{\mathrm{ne}}_{\mathrm{srt}}$ is a solution to
$\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\left[\pi_{\mathrm{u}_{t}}\,G_{t}\,\mathbb{1}\bigl{\{}cG_{t}\leq L_{t}\bigl{\}}\right]=\pi_{0},$ (64)
which is equivalent to
$\sum_{t\in\mathcal{T}}\int_{0}^{L_{t}/c}\pi_{\mathrm{u_{t}}}g_{t}f_{G_{t}}(g_{t})\,\,\mathrm{d}g_{t}=\pi_{0}\frac{T}{\widetilde{T}}$. (65)
Because the right-hand sides of (63) and (65) are equal, it follows that there exists some $\tau\in\mathcal{T}$ such that
$\int_{0}^{L_{\tau}/c}\left(\pi_{\mathrm{u}_{\tau}}+\epsilon\,\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c\,g_{\tau}}{L_{\tau}}\right)\right)g_{\tau}f_{G_{\tau}}(g_{\tau})\,\mathrm{d}g_{\tau}$
$\leq\int_{0}^{L_{\tau}/c}\pi_{\mathrm{u}_{\tau}}g_{\tau}f_{G_{\tau}}(g_{\tau})\,\mathrm{d}g_{\tau}.$ (66)
However, note that
$\left(\pi_{\mathrm{u}_{\tau}}+\epsilon\,\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c\,g_{\tau}}{L_{\tau}}\right)\right)g_{\tau}f_{G_{\tau}}(g_{\tau})\geq\pi_{\mathrm{u}_{\tau}}g_{\tau}f_{G_{\tau}}(g_{\tau}).$ (67)
Combining (66) and (67), we conclude that $\frac{L_{\tau}}{c^{\mathrm{ne}}_{\mathrm{srt}}}\geq\frac{L_{\tau}}{c^{\mathrm{ne}}_{\mathrm{prt}}}$, which implies $c^{\mathrm{ne}}_{\mathrm{prt}}\geq c^{\mathrm{ne}}_{\mathrm{srt}}$.
Proof:
If $\epsilon=0$, then (63) and (65) become identical, which implies $c_{\mathrm{opt}}=c^{\mathrm{ne}}_{\mathrm{prt}}=c^{\mathrm{ne}}_{\mathrm{srt}}$. Meanwhile, for a given $\epsilon\geq 0$, $d_{i}^{\star}(\pi)$ is a solution to:
$\max_{d_{i}\geq 0}\,\,\sum_{t\in\mathcal{T}}\left[\epsilon v_{i}\mathbb{E}\min\{d_{i}G_{t},L_{t}\}-\pi_{\mathrm{u}_{t}}\mathbb{E}(L_{t}-d_{i}G_{t})_{+}\right]-\pi d_{i}{\color[rgb]{0,0,0}.}$
[…]
$\sum_{t\in\mathcal{T}}\left(\pi_{\mathrm{u}_{t}}+\epsilon\tilde{v}_{i}\right)\mathbb{E}\left[G_{t}\mathbb{1}\bigl{\{}d_{i}\left(\frac{\pi_{0}\,T}{\widetilde{T}}\right)G_{t}\leq L_{t}\bigl{\}}\right]=\pi_{0}\frac{T}{\widetilde{T}}$. (68)
[…]
Proof:
The steps follow the same structure as in the proof of Theorem 1. Therefore, we begin by introducing new lemmas analogous to Lemmas 5 and 6.
Lemma 8 (First-order approximation of $c_{\mathrm{prt}}^{\mathrm{ne}}$ around $\epsilon=0$ with heterogeneous periods )
For a sufficiently small $\epsilon>0$, we have
$c_{\mathrm{prt}}^{\mathrm{ne}}=$ $c_{0}-\frac{\epsilon\,\sum_{t\in\mathcal{T}}\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{0}G_{t}}{L_{t}}\right)\,G_{t}\,\mathbb{1}\bigl{\{}c_{0}\,G_{t}\leq L_{t}\bigl{\}}\right]}{\sum_{t\in\mathcal{T}}\pi_{\mathrm{u_{t}}}\mu_{G_{t,\mathrm{tr}}}^{\prime}(c_{0})}$
$+O\left(\epsilon^{2}\right).$ (69)
Lemma 9 (First-order approximation of $c_{\mathrm{cb}}^{\mathrm{ne}}$ around $\epsilon=0$ with heterogeneous periods)
For a sufficiently small $\epsilon>0$, we have
$d_{i}^{\star}\left(\frac{\pi_{0}\,T}{\widetilde{T}}\right)=c_{0}-\sum_{t\in\mathcal{T}}\frac{\epsilon\tilde{v}_{i}\mu_{G_{t,\mathrm{tr}}}(c_{0})}{\pi_{\mathrm{u_{t}}}\mu_{G_{t,\mathrm{tr}}}^{\prime}(c_{0})}+O\left(\epsilon^{2}\right),$ (70)
and therefore,
$c_{\mathrm{cb}}^{\mathrm{ne}}=c_{0}-\sum_{t\in\mathcal{T}}\frac{\epsilon\,\mathbb{E}[\widetilde{V}]\mu_{G_{t,\mathrm{tr}}}(c_{0})}{\pi_{\mathrm{u}}\mu_{G_{t,\mathrm{tr}}}^{\prime}(c_{0})}+O\left(\epsilon^{2}\right).$ (71)
Equipped with these lemmas, we can then proceed by proving
$\sum_{t\in\mathcal{T}}\!\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{0}G_{t}}{L_{t}}\right)\!G_{t}\mathbb{1}\bigl{\{}c_{0}G_{t}\leq L_{t}\bigl{\}}\right]\!<\!\sum_{t\in\mathcal{T}}\!\mathbb{E}[\widetilde{V}]\mu_{G_{t,\mathrm{tr}}}\!(c_{0}).$
We then apply the result from (39) to each operation period; summing the resulting inequalities yields the above result and completes the proof. ∎
Appendix E Proof of Lemma 8 and Lemma 9
E-A Proof of Lemma 8
We can rewrite (62) as:
$\underbrace{\pi_{0}\frac{T}{\widetilde{T}}-\sum_{t\in\mathcal{T}}\left[\pi_{\mathrm{u}_{t}}\,\mu_{G_{t,\mathrm{tr}}}(c^{\mathrm{ne}}_{\mathrm{prt}})-\epsilon\,k_{t}\left(c^{\mathrm{ne}}_{\mathrm{prt}}\right)\right]}_{:=\phi_{\mathrm{prt}}\left(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}}\right)}=0.$ (72)
Here, we overload our notation by letting $\phi_{\mathrm{prt}}\left(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}}\right)$ denote the left-hand side of the equation and defining
$k_{t}(c^{\mathrm{ne}}_{\mathrm{prt}}):=\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{\mathrm{prt}}^{\mathrm{ne}}G_{t}}{L_{t}}\right)\,G_{t}\,\mathbb{1}\left\{c^{\mathrm{ne}}_{\mathrm{prt}}G_{t}\leq L_{t}\right\}\right],$
[…]
E-B Proof of Lemma 9
We begin by rewriting (68) in the following form:
$\underbrace{\pi_{0}\,\frac{T}{\widetilde{T}}-\sum_{t\in\mathcal{T}}\left[\left(\pi_{\mathrm{u}_{t}}+\epsilon\,\tilde{v}_{i}\right)\mu_{G_{t,\mathrm{tr}}}\left(d_{i}\left(\frac{\pi_{0}T}{\widetilde{T}}\right)\right)\right]}_{:=\phi_{\mathrm{cb},i}\left(\epsilon,d_{i}\left(\frac{\pi_{0}T}{\widetilde{T}}\right)\right)}=0,$ (73)
[…]

### v2

@@
-Market-Driven Equilibria for Distributed Solar Panel Investment
+Extended Version: Characterizing Distributed Photovoltaic Panel Investment Equilibria
 Abstract
-This study investigates market-driven long-term investment decisions in distributed solar panels by individual investors.
-We consider a setting where investment decisions are driven by expected revenue from participating in short-term electricity markets over the panel’s lifespan.
+This study investigates long-term investment decisions in distributed photovoltaic panels by individual investors.
+We consider a setting where investment decisions are driven by expected revenue from participating in short-term electricity markets over the panel lifespan.
 These revenues depend on short-term markets equilibria, i.e., prices and allocations, which are influenced by aggregate invested panel capacity participating in the markets.
@@
 We also establish that the contract-based market leads to over-investment when the extra valuations of users for solar energy are small.
-Finally, we validate our theoretical findings through numerical experiments.
+Finally, we validate our theoretical results through numerical experiments.
 I Introduction
-The increasing adoption of renewable energy, particularly solar photovoltaic systems, plays an important role in modern power systems.
+The increasing adoption of renewable energy, particularly photovoltaic (PV) panels, plays an important role in modern power systems.
 While utility-scale solar projects contribute significantly to overall capacity, distributed rooftop installations are playing an increasingly prominent role in reshaping the electricity landscape.
-Notably, as of 2022, approximately 140 GW of solar capacity had been installed in the U.S., with more than 35% of it coming from installations by households and commercial entities [1].
-Unlike centralized projects planned and operated by utilities, distributed solar investments are driven by decisions of individual households and businesses.
+Notably, by the end of 2024, the U.S.
+had installed about 175 GW of solar capacity, over 31% of which came from residential and commercial installations [1].
+Unlike centralized projects planned and operated by utilities, distributed solar investments at the distribution level reflect individual adoption and investment decisions by households and businesses.
 Several factors influence these decisions.
-A key factor is the financial return from the investment, driven by the expected revenue earned in short-term distribution-level markets over the panels’ lifespan, where that revenue is determined by the markets’ equilibria.
+A key factor is the financial return from the investment, driven by the expected revenue earned in short-term distribution-level markets over the panel lifespan, where that revenue is determined by the markets’ equilibria.
 Meanwhile, these short-term equilibria depend on the aggregate invested panel capacity available in the markets, which is shaped by the long-term investment equilibrium that emerges from interactions among individual investors.
@@
 A precise characterization of this relationship is crucial for understanding the emergence of investment equilibria.
-This insight is not only important for utilities and independent system operators, who should account for distributed capacity growth in their planning studies, but also for policymakers and market designers aiming to understand the long-term impact of diverse market designs on the solar panel adoption.
-This motivates us to study the following critical open question: How do short-term market equilibria, arising from diverse solar market mechanisms, shape the corresponding long-term distributed investment equilibrium?
+This insight is not only important for utilities and independent system operators, who should account for distributed capacity growth in their planning studies, but also for policymakers and market designers aiming to understand the long-term impact of diverse market designs on PV panel adoption.
+This motivates us to study the following critical open question [^1] We note that a few studies in related settings (e.g., conventional generation investment) also link long-run investment equilibria to short-run operational equilibria.
+However, they largely focus on spot markets [2, 3] and either do not yield closed-form characterizations [2] or incorporate detailed consumer modeling (e.g., heterogeneity) [2, 3].: How do short-term market equilibria, arising from diverse solar market mechanisms, shape the corresponding long-term distributed investment equilibrium?
 I-A Organization and Contributions
-To address this problem, we consider a large population of individual investors and aim to characterize the long-term investment equilibrium that emerges from the previously described feedback loop.
-We first develop an investment model, formulated as a non-atomic game, that links long-term investment decisions to expected revenues under a given market design (Section II).
-To characterize these revenues, we introduce a unified modeling framework for short-term markets (Section III).
-This framework allows us to formally define and compare different market designs by capturing key features such as product differentiation and ex-ante/ex-post trading.
-We then apply it to three representative mechanisms that differ in these key features (Section IV): (a) a single-product real-time market, (b) a product-differentiated real-time market, and (c) a contract-based panel market.
-For each, we first derive analytical characterizations of the short-term market equilibria and then use the investment model to derive the associated long-term equilibrium aggregate capacities (Section V).
+To address this open problem, we consider a large population of individual investors and aim to characterize the long-term investment equilibrium that emerges from the previously described feedback loop.
+We first develop an investment model, formulated as a non-atomic game, where each player is infinitesimal with negligible market power, reflecting the large number of small investors in distribution-level settings.
+This model links long-term investment decisions to expected revenues under a given market design (Section II).
+To characterize these revenues, we introduce a unified modeling framework for different types of short-term markets (Section III) that allows us to formally define and compare different market designs.
+Our unified market model captures key features such as product differentiation and ex-ante/ex-post trading.
+We then apply it to three representative mechanisms that differ in these features (Section IV).
+First, we study (a) a single-product real-time market similar to today’s real-time electricity markets; then (b) a new product-differentiated real-time market to assess the impact of distinguishing clean solar energy from conventional grid energy; and finally (c) a new contract-based panel market that operates on a much slower timescale than real-time markets.
+For each, we first derive analytical characterizations of the short-term market equilibria and then use the investment model to obtain and compare the associated long-term equilibrium aggregate capacities (Section V).
 Our analysis, initially assuming homogeneous operation periods, is extended to handle heterogeneous operation periods (Section VI), followed by numerical experiments (Section VII).
@@
   1. (a)
-    The non-atomic game-based investment model encapsulates how individual investors impact each other through their aggregate investment decisions, offering an analytically tractable way to link short-term market equilibria to long-term aggregate investment equilibrium.
+    We offer an analytically tractable way to link short-term market equilibria to long-term aggregate investment equilibrium through a non-atomic game-based investment model that captures how individual investors affect one another through aggregate investment decisions.
+This model can also represent common pricing schemes (e.g., feed-in tariffs and time-of-use rates) as long-run averages of short-term market prices that co-evolve with investment decisions, rather than exogenous fixed tariffs.
   2. (b)
-    The proposed unified modeling structure provides closed-form characterizations of short-term market equilibria across diverse mechanism designs, accommodating both ex-ante and ex-post trading as well as heterogeneous consumer preferences.
+    We develop a unified market-modeling framework that delivers closed-form characterizations of short-term market equilibria across diverse market mechanisms, accommodating both ex-ante and ex-post trading as well as heterogeneous consumer preferences.
   3. (c)
-    These analytical characterizations enable us to theoretically establish that the single-product real-time market, where solar energy is pooled with conventional grid energy and traded ex-post, leads to under-investment.
-In contrast, the product-differentiated market, where solar energy is traded ex-post and may be sold at a premium due to heterogeneous user preferences, yields the socially optimal capacity.
-Finally, the contract-based market, where ex-ante panel capacity is traded instead of realized generation, can result in over-investment when users’ extra valuations for solar are low.
+    These analytical characterizations yield new insights into how different market mechanisms shape PV investment.
+Specifically, we theoretically establish that the single-product real-time market, where solar energy is pooled with conventional grid energy and traded ex-post, leads to under-investment.
+Thus, we believe there is significant value in considering alternative market mechanisms.
+In particular, the product-differentiated market, where solar energy is traded ex-post and may be sold at a premium due to heterogeneous user preferences, yields the socially optimal capacity.
+Finally, the contract-based market, where ex-ante panel capacity is traded instead of realized generation, can result in over-investment when users’ extra valuations for solar energy are small.
+Such over-investment can be desirable under economies of scale, where higher aggregate deployment lowers the cost of PV investment.
 I-B Related Literature
 This work is related to the intersection of two broad lines of literature: distribution-level electricity market mechanisms and long-term investment in solar panels.
-Most existing studies on the investment side of the literature adopt a centralized perspective, focusing on the optimal strategy of a single decision-maker such as a utility planner or an individual investor (see, e.g., [2, 3, 4, 5, 6]).
+Most existing studies on the investment side of the literature adopt a centralized perspective, focusing on the optimal strategy of a single decision-maker such as a utility planner or an individual investor (see, e.g., [4, 5, 6]).
 However, this line of work neglects the interdependencies of investment decisions in distributed settings, where the actions of one investor can influence the returns of others through shared market mechanisms.
@@
 In contrast, service contract-based mechanisms typically offer less volatile prices in exchange for a fixed level of electricity service, relying on ex-ante statistics such as those related to random solar generation.
-Extensive research has explored the relative merits of these retail market designs, examining factors such as economic efficiency [7, 8, 9, 10], price and bill volatility [11, 12, 13], and environmental impacts [14].
-In addition, some studies have examined another facet of market design: whether electricity from solar and conventional sources should be treated as differentiated products in electricity markets, given their distinct operational characteristics and emissions [15, 16, 17, 18, 19].
-Despite the contributions of all these reviewed studies, they all focus primarily on short-term performance criteria and overlook the long-term impacts of their designed markets on investment decisions.
-The most closely related paper to ours is [20], which analyzes the equilibrium panel investment capacities under feed-in-tariff and sharing economy markets.
-Nonetheless, they adopt a finite-player game to analyze the setup involving large investors.
-In contrast, we focus on the setting with a large number of small investors.
-Through a non-atomic game formulation, we can analytically compare equilibrium capacities driven by various market mechanisms, for which they have to resort to numerical simulation.
+Extensive research has explored the relative merits of these retail market designs, examining factors such as economic efficiency [7, 8], price and bill volatility [9, 10], and environmental impacts [11].
+In addition, some studies have examined another facet of market design: whether electricity from solar and conventional sources should be treated as differentiated products in electricity markets, given their distinct operational characteristics and emissions [12, 13, 14].
+Despite the contributions of these studies[7, 8, 9, 10, 11, 12, 13, 14], they all focus primarily on short-term performance criteria and overlook the long-term impacts of their designed markets on investment decisions.
+The most closely related paper to ours is [15], which studies equilibrium PV investment capacities under feed-in-tariff and sharing-economy markets.
+However, it adopts a finite-player game with large investors, whereas we consider a large population of small investors.
+Our non-atomic game formulation enables analytical comparisons of equilibrium capacities driven by different market mechanisms, for which they have to resort to numerical simulations.
+While non-atomic game formulations have been used in other contexts (e.g., workplace EV charging pricing [16] and distributed storage investment [17]), they have not been used in an investment framework that links short-term market equilibria under diverse market designs to long-term investment equilibrium.
 This paper generalizes our previous conference work [18] by (a) extending our theoretical results to setups with heterogeneous operation periods, (b) presenting rigorous mathematical proofs, and (c) demonstrating our theoretical results by performing numerical experiments.
 II Investment Model
-A large group of potential investors considers installing solar panels and becoming panel owners.
-Their decisions weigh upfront capital and installation costs against expected revenues from participating in short-term electricity markets over the panels’ lifespan.
+A large group of potential investors considers installing PV panels and becoming panel owners.
+Their decisions weigh upfront capital and installation costs against expected revenues from participating in short-term electricity markets over the panel lifespan.
 An overview of the setup is shown in Fig.
@@
 II-A Connecting Planning and Operation Timescales
-Investors typically consider a 25-year lifespan for solar panels [22].
-However, estimating expected revenue at the time of investment is challenging, as reliable long-term forecasts of solar output and market conditions are rarely available.
-A common approach is to use a shorter planning period (e.g., one year), estimate revenues from historical data, and scale [^1] This step may include interest rate adjustments and net present value computations.
-the result by the number of such intervals over the panel’s lifespan.
+Investors typically assume a 25-year PV panel lifespan [19].
+Accordingly, at the investment stage, estimating expected revenue over such a long horizon is challenging because long-term forecasts of solar generation and market conditions are rarely available.
+In response, a common approach is to use a shorter planning period (e.g., one year), estimate revenues from historical data, and scale [^2] This step may include interest rate adjustments and net present value computations.
+the result by the number of such intervals over the panel lifespan.
 The appropriate length of the planning period should be selected based on the availability and granularity of historical data.
-When the planning period is long, there may be multiple operation periods within it.
+Since the planning period is long, it contains multiple operation periods, each on a shorter timescale (e.g., an hour or a day).
 We consider a finite-horizon discrete-time planning window consisting of operation periods $\mathcal{T}=\{1,\dots,T\}$, indexed by $t$, where $T$ is the total number of such periods.
@@
 Thus, the expected revenue remains consistent across operation periods.
-We estimate it for a single representative period (e.g., as shown in Fig.
-1), then scale it up to the planning horizon and ultimately to the full lifespan.
-The total expected revenue is therefore obtained by multiplying the representative period’s revenue by a scaling factor $\widetilde{T}$, which captures the number of such periods over the panel’s lifespan and incorporates any interest-rate adjustments.
+Therefore, we can first estimate this revenue over a single operation period, which we refer to as the representative operation period as shown in Fig.
+1, and then scale it to obtain the expected revenue for the planning horizon and, ultimately, for the panel lifespan.
+The total expected revenue is therefore obtained by multiplying the representative operation period’s revenue by a scaling factor $\widetilde{T}$, which captures the number of all operation periods over the panel lifespan and incorporates any interest-rate adjustments.
 Remark 1 (More granular operation model)
@@
 II-B Panel Investment Game
-Considering a large number of small investors each of whom does not have market power, we model the group of investors as the continuous interval $\mathcal{I}_{\mathrm{inv}}:=[0,1]$.
-We consider the case where each investor $i\in\mathcal{I}_{\mathrm{inv}}$ decides whether to install a solar panel with a fixed capacity $\overline{c}$.
+Considering a large number of small investors, each of whom does not have market power and is infinitesimal relative to the market (so individual actions do not affect outcomes while the aggregate does), we model their competition as a non-atomic game [20].
+Consequently, we represent the investors population as the continuum $\mathcal{I}_{\mathrm{inv}}:=[0,1]$.
+We consider the case where each investor $i\in\mathcal{I}_{\mathrm{inv}}$ decides whether to install a PV panel with a fixed capacity $\overline{c}$.
 Let the decision of investor $i$ be $x_{i}\in\{0,1\}$.
 The total panel capacity that the investors will install is then
-$c=\overline{c}\int_{\mathcal{I}_{\mathrm{inv}}}x_{i}\,\,\mathrm{d}i.$ (1)
+$c=\overline{c}\int_{\mathcal{I}_{\mathrm{inv}}}x_{i}\,\,\mathrm{d}i.\vskip-2.5pt$ (1)
 Let the capital and installation cost for a unit panel capacity be $\pi_{0}$. We assume that the panel upfront cost is attractive, at least when its solar production can be sold at the utility rate:
-${\color[rgb]{0,0,0}c}\pi_{0}\leq\widetilde{T}\pi_{\mathrm{u}}\mathbb{E}[{\color[rgb]{0,0,0}c}G],$ (2)
-where $\pi_{\mathrm{u}}>0$ is the fixed price at which the utility provides backstop electricity to buyers, and $G$ denotes solar generation per unit panel capacity, typically measured in $\mathrm{W}/\mathrm{m}^{2}$ or $\mathrm{GW}/\mathrm{km}^{2}$.
+${\color[rgb]{0,0,0}c}\pi_{0}\leq\widetilde{T}\pi_{\mathrm{u}}\mathbb{E}[{\color[rgb]{0,0,0}c}G],\vskip-1.99997pt$ (2)
+where $\pi_{\mathrm{u}}>0$ is the fixed price at which the utility provides electricity to buyers, and $G$ denotes solar generation per unit panel capacity, typically measured in $\mathrm{W}/\mathrm{m}^{2}$ or $\mathrm{GW}/\mathrm{km}^{2}$.
 This solar generation is modeled as a random variable with a cumulative distribution function $F_{G}$ and a probability density function $f_{G}$.
@@
 Thus, the payoff for investor $i$ from installing panel capacity $\overline{c}$ and selling the generated electricity in a short-term market $m$ (where $m$ serves as an index for different short-term market mechanisms) is given by the expected total revenue minus the investment cost, expressed as:
-$\Pi^{\mathrm{inv}}_{m,i}(x_{i},c)=\left[(\Pi_{m}^{\mathrm{s}}(c)/c)-\pi_{0}\right]\overline{c}\,x_{i},$ (3)
-where $\Pi_{m}^{\mathrm{s}}(c)$ represents the expected total revenue received by solar panel owners under the short-term market $m$.
-Given the players $\mathcal{I}_{\mathrm{inv}}$, actions $x=\{x_{i}\}_{i\in\mathcal{I}_{\mathrm{inv}}}$, and payoff (3), we have defined the solar panel investment game, which is an aggregate game where the dependence of investor $i$’s payoff on the other investors’ decisions is only through the total panel capacity $c$.
+$\Pi^{\mathrm{inv}}_{m,i}(x_{i},c)=\left[(\Pi_{m}^{\mathrm{s}}(c)/c)-\pi_{0}\right]\overline{c}\,x_{i}.$ (3)
+Here, $\Pi_{m}^{\mathrm{s}}(c)$ represents the expected total revenue received by panel owners under the short-term market $m$, to be defined individually for each market in Section III.
+Each investor $i\in\mathcal{I}_{\mathrm{inv}}$ chooses its investment decision to maximize $\Pi^{\mathrm{inv}}_{m,i}(x_{i},c)$.
+Given the players $\mathcal{I}_{\mathrm{inv}}$, actions $x=\{x_{i}\}_{i\in\mathcal{I}_{\mathrm{inv}}}$, and payoff (3), we have defined the panel investment game, which is an aggregate game where investor $i$’s payoff depends on the other investors’ decisions only through the total panel capacity $c$.
 A Nash equilibrium (NE) for this game is defined as:
 Definition 1 (NE for panel investment game)
-For a market mechanism $m$, the aggregate panel capacity $c$ constitutes an NE for the solar panel investment game with market structure $m$ if there exists a collection of investment decisions $x$ such that $\Pi_{m,i}^{\mathrm{inv}}(x_{i},c)\geq\Pi_{m,i}^{\mathrm{inv}}(x_{i}^{\prime},c)$, for all $x_{i}^{\prime}\in\{0,1\}$ and $i\in\mathcal{I}_{\mathrm{inv}}$, and (1) holds.
+The aggregate panel capacity $c$ constitutes an NE for the panel investment game under market mechanism $m$ if there exists a collection of investment decisions $x$ such that $\Pi_{m,i}^{\mathrm{inv}}(x_{i},c)\geq\Pi_{m,i}^{\mathrm{inv}}(x_{i}^{\prime},c)$, for all $x_{i}^{\prime}\in\{0,1\}$ and $i\in\mathcal{I}_{\mathrm{inv}}$, and (1) holds.
 Under this condition, no investor will have an incentive to unilaterally deviate from the NE decision given the aggregate capacity that is consistent with everyone’s investment decision.
@@
 The market consists of buyers (i.e., electricity consumers), sellers [^3] We acknowledge the possibility that electricity prosumers can be both buyers and sellers at the same time.
-For the market mechanisms that we will consider in this paper, a prosumer can be modeled as the superposition of a buyer and a seller.
+For the market mechanisms that we will consider in this paper, a prosumer can be modeled as the superposition of a buyer and a seller since buying and selling prices are uniform within each market.
 This observation may not apply to other markets where buying and selling prices are non-uniform.
-(i.e., panel owners), and a utility company that provides backstop supply when solar generation alone cannot fully satisfy demand.
-Similar to the investment game, we model the groups of a large number of buyers and sellers as continuous intervals.
+(i.e., the collection of investors who choose to invest), and a utility company that provides supply when solar generation alone cannot fully satisfy demand.
+Similar to the investment game, we model the groups of a large number of buyers and sellers as continuous intervals through a non-atomic game formulation.
+[^4] Non-atomic game formulations [20] differ from non-atomistic market models [21], where some players are large enough to exercise market power.
 Define the set of buyers by the interval $\mathcal{I}_{\mathrm{b}}:=[0,1]$.
 In our analysis, we consider a situation where environmentally conscious electricity consumers have an extra heterogeneous valuation for solar energy over the conventional energy supplied by the utility.
-For a buyer $i\in\mathcal{I}_{\mathrm{b}}$, we denote their extra valuation for each unit of load served by solar as $v_{i}\in\mathbb{R}_{+}$, referred to as the solar premium.
-The distribution of $v_{i}$ can be characterized by a function $F_{V}$ defined as $F_{V}(v)=\int_{i\in\mathcal{I}_{\mathrm{b}}}\,\mathbb{1}\{v_{i}\leq v\}\,\,\mathrm{d}i$, where $\mathbb{1}$ is the indicator function.
+For a buyer $i\in\mathcal{I}_{\mathrm{b}}$, we denote their fixed extra valuation for each unit of load served by solar as $v_{i}\in\mathbb{R}_{+}$, referred to as the solar premium.
+Accordingly, their willingness to pay for a unit of solar-served load is $\pi_{\mathrm{u}}+v_{i}$.
+Heterogeneity is captured by the distribution of $\{v_{i}\}_{i\in\mathcal{I}_{\mathrm{b}}}$ across buyers, characterized by a function $F_{V}$ defined as $F_{V}(v)=\int_{i\in\mathcal{I}_{\mathrm{b}}}\,\mathbb{1}\{v_{i}\leq v\}\,\,\mathrm{d}i$, where $\mathbb{1}$ is the indicator function.
 For simplicity, we assume that $F_{V}$ is differentiable and strictly increasing on $0,\overline{v}]$, where $\overline{v}\geq 0$.
@@
 Each seller $i\in\mathcal{I}_{\mathrm{s}}$ can sell any quantity of solar generation within the set $\mathcal{Q}_{i}^{\mathrm{s}}=[0,cG]$, which depends on the realized value of $G$.
-The seller payoff function is
-$\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)=\pi q^{\mathrm{s}}_{i},$
+The seller payoff function for the representative operation period is
+$\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)=\pi q^{\mathrm{s}}_{i},$ (4)
 where market equilibrium price $\pi$ and cleared quantity $q^{\mathrm{s}}_{i}$ will be defined momentarily.
 Given the inelastic load $L$, a buyer $i\in\mathcal{I}_{\mathrm{b}}$ can purchase any quantity of solar within the range $\mathcal{Q}^{\mathrm{b}}_{i}=[0,L]$.
-Any remaining demand not met by solar will be fulfilled by the utility backstop at the price $\pi_{\mathrm{u}}$.
-The buyer payoff function is
-$\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)=(v_{i}-\pi)q^{\mathrm{b}}_{i}-\pi_{\mathrm{u}}(L-q^{\mathrm{b}}_{i}),$
-where $q_{i}^{\mathrm{b}}$ is the cleared quantity for buyer $i$.
+Any remaining demand not met by solar will be fulfilled by the utility supply at the price $\pi_{\mathrm{u}}$.
+The buyer payoff function for the representative operation period is
+$\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)=(v_{i}-\pi)q^{\mathrm{b}}_{i}-\pi_{\mathrm{u}}(L-q^{\mathrm{b}}_{i}),$ (5)
+where $q_{i}^{\mathrm{b}}$ is the cleared quantity for the buyer.
+The first term in (5) is the buyer’s surplus in the market, and the second term accounts for the payment for the residual demand served by the utility.
 Denote $q^{\mathrm{s}}:=\{q^{\mathrm{s}}_{i}\}_{i\in\mathcal{I}_{\mathrm{s}}}$ and $q^{\mathrm{b}}:=\{q^{\mathrm{b}}_{i}\}_{i\in\mathcal{I}_{\mathrm{b}}}$.
@@
 Such outcomes can arise, for instance, from buyers and sellers submitting bids to a market operator who clears the market.
-In our non-atomic game model, where all agents are infinitesimally small and act as price takers, it can be shown that an NE of such a bidding process corresponds to a CE that we consider here.
-Detailed models of price formation have been extensively studied in prior work (see references in Section I) and are beyond the scope of this paper.
-With the equilibrium price $\pi$ and the cleared quantities $q^{\mathrm{s}}_{i}$ for sellers in the product-differentiated real-time market, the total expected revenue for sellers with an aggregate panel capacity $c$ over the panels’ lifespan can be calculated as
+In our non-atomic game model, where all agents are infinitesimally small and act as price takers, we show that an NE of such a bidding process corresponds to a CE that we consider here (see Appendix A).
+With the equilibrium price $\pi$ and the cleared quantities $q^{\mathrm{s}}_{i}$ for sellers in the product-differentiated real-time market, the total expected revenue for sellers with an aggregate panel capacity $c$ over the panel lifespan can be calculated as
 $\Pi^{\mathrm{s}}_{\mathrm{prt}}(c)=\widetilde{T}\,\mathbb{E}\left[\int_{\mathcal{I}_{\mathrm{s}}}\Pi^{\mathrm{s}}_{i}(q_{i}^{\mathrm{s}},\pi)\,\,\mathrm{d}i\right].$ (6)
@@
 Nevertheless, for notational consistency, we will reuse the notation from the previous subsection to describe outcomes in the contract-based setting.
+Remark 3 (Time-scale of contract-based market)
+In this section, we assume contracts cover only a single operation period.
+Although contracts are often long-term in practice, assuming operation periods are identically distributed implies that market outcomes repeat across periods.
+Thus, contracting for one representative period yields the same outcomes as contracting for multi-period blocks.
+Importantly, contracts are still set ex ante.
+Section VI relaxes this assumption and considers heterogeneous periods and multi-period agreements.
 Since seller $i\in\mathcal{I}_{\mathrm{s}}$ can rent out any portion of their panel capacity $c$, the set of feasible amount s to sell is $\mathcal{Q}_{i}^{\mathrm{s}}=[0,c]$.
@@
 where $(z)_{+}:=\max(z,0)$, and $q_{i}^{\mathrm{s}}$ and $q_{i}^{\mathrm{b}}$ are the cleared quantities (i.e., panel capacities) for the seller and buyer, respectively.
+In (7), the first two terms capture the buyer’s expected surplus in the market, while the last term is the expected payment to the utility for any residual demand.
 Adapting the same CE concept as in Definition 2, with the traded product and payoff functions updated, and given the equilibrium price $\pi$ and seller cleared quantity $q^{\mathrm{s}}$, we can express the total sellers revenue for the contract-based market as
@@
 Remark 4 (Unified modeling framework)
-The proposed market model serves as a general framework capable of representing a variety of solar market mechanisms, provided that the traded product and the payoff structures for buyers and sellers are appropriately defined.
+The proposed market model serves as a general framework capable of representing a variety of solar market mechanisms, provided that the traded product and the payoff structures for buyers and sellers are appropriately defined such that a CE exists.
 For example, both real-time pricing and contract-based arrangements can be formulated within this framework, with key distinctions captured through the specification of the traded product, the feasible sets of buyers and sellers, and the corresponding payoff functions.
 We abstract away additional implementation-specific features, as our primary objective is to investigate how different market structures shape long-term investment outcomes.
-While these mechanisms may differ in operational frequency, often leading to variations in price volatility as discussed in prior work [11, 12, 13], such factors do not directly affect the investment equilibrium when only expected returns matter.
+While these mechanisms may differ in operational frequency, often leading to variations in price volatility as discussed in prior work [9, 10], such factors do not directly affect the investment equilibrium when only expected returns matter.
 Now that we have completed the modeling of these market mechanisms, we proceed to analyze how the corresponding CE can be derived for each.
-IV Competitive Equilibria of Solar Markets
-In this section, we derive the expected revenue for panel owners as a function of panel capacity $c$.
+IV Competitive Equilibria of Short-Term Markets
+In this section, we derive the expected revenue for panel owners as a function of panel capacity $c$. Proofs are provided in the appendices.
 Lemma 1 (CE of real-time market mechanisms)
 A CE for product-differentiated real-time market (prt) and single-product real-time market (srt), are listed in Table I, where $\overline{F}_{V}(\cdot):=1-F_{V}(\cdot)$ is the complementary cumulative distribution function of $v_{i}$, and $\overline{\pi}^{\mathrm{b}}_{i}:=\pi_{\mathrm{u}}+v_{i}$.
-Furthermore, the expected total revenue for solar panel owners under these two market mechanisms is
+Furthermore, the expected total revenue for panel owners under these two market mechanisms is
 $\Pi^{\mathrm{s}}_{\mathrm{prt}}(c)=\widetilde{T}\,\mathbb{E}\left[\left(\pi_{\mathrm{u}}+\overline{F}^{-1}_{V}\left(\frac{cG}{L}\right)\right)\,cG\,\mathbb{1}\bigl\{cG\leq L\bigl\}\right],$ (9)
@@
 Market $\pi$ $q_{i}^{\mathrm{s}}$ $q_{i}^{\mathrm{b}}$
-Abundant supply
-$cG>L$
-prt 0 $L$ $L$
+Abundant supply $cG>L$ prt 0 $L$ $L$
 srt 0 $L$ $L$
-Limited supply
-$cG\leq L$
-prt $\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)$ cGcG $L\mathbb{1}\{\overline{\pi}^{\mathrm{b}}_{i}\geq\pi\}$
+Limited supply $cG\leq L$ prt $\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)$ cGcG $L\mathbb{1}\{\overline{\pi}^{\mathrm{b}}_{i}\geq\pi\}$
 srt $\pi_{\mathrm{u}}$ cGcG cGcG
@@
 When the solar supply is insufficient to cover the load, the market clearing outcomes of the two market types become different.
-Since the utility backstop is required in this case, the market price for the single-product real-time market will be $\pi_{\mathrm{u}}$.
+Since the utility supply is required in this case, the market price for the single-product real-time market will be $\pi_{\mathrm{u}}$.
 When solar is traded as a separate product, however, we see that the scarce solar production $cG$ is allocated among the buyers with higher solar premiums first.
 The solar market price turns out to have a premium over the utility-supplied energy, driven by the competition among the buyers.
+Accordingly, substituting $\pi$ and $q_{i}^{\mathrm{s}}$ from Table I into (4) yields individual seller payoffs; aggregating via (6) then gives (9).
+The expression in (10) follows by a similar calculation under the single-product market model.
 We now proceed to characterize the CE for the contract-based market.
@@
 $q_{i}^{\mathrm{s}}=c,\quad i\in\mathcal{I}_{\mathrm{s}};\quad q_{i}^{\mathrm{b}}=d_{i}^{\star}(\pi),\quad i\in\mathcal{I}_{\mathrm{b}},$
-and $\pi$ is the solution to $c=\widehat{d}^{\star}(\pi).$ Moreover, the total revenue for solar panel owners under the contract-based market is
+and $\pi$ is the solution to $c=\widehat{d}^{\star}(\pi).$ Moreover, the total revenue for panel owners under the contract-based market is
 $\Pi^{\mathrm{s}}_{\mathrm{cb}}(c)=\widetilde{T}c\pi.$ (14)
 V Equilibrium Panel Capacities
-Building on the short-term market outcomes derived in the previous section, we next study the investment model to characterize the resulting equilibrium panel investment capacities.
+Given the short-term market outcomes derived in the previous section, we next apply the investment model to characterize the resulting equilibrium panel investment capacities.
 Lemma 4 (NE panel capacities)
@@
 When assumption (2) is violated, the stated equilibrium condition no longer applies, and $c_{\mathrm{srt}}^{\mathrm{ne}}=0$.
-While Lemma 4 enables us to use numerical simulations to compare the equilibrium capacities for each market mechanism, deeper insight comes from a theoretical comparison of the resulting aggregate capacities.
+While Lemma 4 enables us to compare the equilibrium capacities across market mechanisms via numerical simulations, deeper insight comes from a theoretical comparison of the resulting aggregate capacities.
 To enable a more meaningful comparison, we first define a benchmark capacity as follows.
@@
 The social welfare optimal panel capacity is then a solution to the following optimization problem
-$\max_{c\in\mathbb{R}_{+}}\enspace$ $\widetilde{T}\,\mathbb{E}\left[v^{\star}(c,G)L-\pi_{\mathrm{u}}(L-cG)_{+}\right]-\pi_{0}c.$ (18)
+$\max_{c\in\mathbb{R}_{+}}\kern 5.0pt$ $\widetilde{T}\,\mathbb{E}\left[v^{\star}(c,G)L-\pi_{\mathrm{u}}(L-cG)_{+}\right]-\pi_{0}c,$ (20)
+where the objective function of (20) sums consumers and investors payoffs; internal monetary transfers between them cancel out.
 The following result provides an analytical characterization of this capacity.
@@
     General case: For any $\epsilon\geq 0$,
-    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}.$ (20)
+    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}.\vskip-2.5pt$ (22)
   2. 2.
     Small solar premium: If $\epsilon$ is in a neighborhood of 0 and $c^{\mathrm{ne}}_{\mathrm{srt}}$ is large such that
-    $(1-\delta)r_{0}\leq f_{G}(g)\leq(1+\delta)r_{0},$ (21)
+    $(1-\delta)r_{0}\leq f_{G}(g)\leq(1+\delta)r_{0},\vskip-1.49994pt$ (23)
     holds for all $g\leq L/c^{\mathrm{ne}}_{\mathrm{srt}}$ with some $r_{0}\in\mathbb{R}_{+}$ and $\delta>0$ sufficiently small, then
-    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}\lesssim c^{\mathrm{ne}}_{\mathrm{cb}},$ (22)
+    $c^{\mathrm{ne}}_{\mathrm{srt}}\leq c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}\lesssim c^{\mathrm{ne}}_{\mathrm{cb}},\vskip-2.5pt$ (24)
     where the last inequality holds in the sense that
-    $c^{\mathrm{ne}}_{\mathrm{cb}}-c^{\mathrm{ne}}_{\mathrm{prt}}\geq{\color[rgb]{0,0,0}\beta\epsilon}-O\left(\epsilon^{2}\right),$ (23)
+    $c^{\mathrm{ne}}_{\mathrm{cb}}-c^{\mathrm{ne}}_{\mathrm{prt}}\geq{\color[rgb]{0,0,0}\beta\epsilon}-O\left(\epsilon^{2}\right),\vskip-2.5pt$ (25)
     with $\beta$ being a positive constant.
@@
     $c^{\mathrm{ne}}_{\mathrm{srt}}=c^{\mathrm{ne}}_{\mathrm{prt}}=c_{\mathrm{opt}}=c^{\mathrm{ne}}_{\mathrm{cb}}.$ (26)
-The first interesting result from (20) in Theorem 1 is that the single-product real-time market, in general, leads to an under-investment in solar panels relative to social welfare optimal panel capacity.
-In fact, with $\epsilon>0$, it is easy to identify a distribution $F_{\widetilde{V}}$ such that $c^{\mathrm{ne}}_{\mathrm{srt}}<c_{\mathrm{opt}}$, and therefore the single-product real-time energy market leads to under-investment.
+The first interesting result from (22) in Theorem 1 is that the single-product real-time market, in general, leads to an under-investment in panels relative to social welfare optimal panel capacity.
+In fact, with $\epsilon>0$, it is easy to identify a distribution $F_{\widetilde{V}}$ such that $c^{\mathrm{ne}}_{\mathrm{srt}}<c_{\mathrm{opt}}$.
 While one could argue that the single-product real-time market is easier to implement, particularly given the limitations of current metering infrastructure, it results in a loss of social welfare when viewed through the lens of long-term investment equilibrium.
-The second observation form (20) is that the product-differentiated real-time market is guaranteed to achieve the socially optimal panel investment.
+The second observation [^6] In this case, an analytical comparison between $c_{\mathrm{prt}}^{\mathrm{ne}}$ and $c_{\mathrm{cb}}^{\mathrm{ne}}$ is not available; nevertheless, our numerical results indicate that $c_{\mathrm{cb}}^{\mathrm{ne}}>c_{\mathrm{prt}}^{\mathrm{ne}}$ in most realistic setups.
+from (22) is that the product-differentiated real-time market is guaranteed to achieve the socially optimal panel investment.
 This stems from the fact that the product-differentiated market indeed allocates solar energy in the same way as the social welfare optimal allocation in (19).
-This leads to the same expected revenue for solar panel owners and therefore the same condition for the resulting aggregate panel capacities.
+This leads to the same expected revenue for panel owners and therefore the same condition for the resulting aggregate panel capacities.
+Nonetheless, implementing such a market in practice is nontrivial and would require granular metering, verification, and settlement to track solar deliveries.
 When the electricity users are unwilling to pay much more for solar energy (i.e., when $\epsilon$ is small), we establish a qualitative relation between the contract-based market equilibrium capacity, $c^{\mathrm{ne}}_{\mathrm{cb}}$, and the social welfare optimal capacity, $c_{\mathrm{opt}}$.
@@
 So far, we have assumed that the values/distributions of key parameters, such as loads and solar irradiation, are identical across operation periods.
-In practice, however, these values/distributions may vary over time due to factors such as seasonality, time of day, and day of the week.
+In practice, however, these values/distributions may vary over time (e.g., by season, time of day, and day of week).
 In this section, we extend our analysis to incorporate this temporal heterogeneity.
-In our extended model, each period $t$ is characterized by a distinct set of parameters: the total consumer load $L_{t}$, the solar generation per unit panel $G_{t}$ with associated probability density function $f_{G_{t}}$, and the utility backstop price $\pi_{\mathrm{u}_{t}}$.
-Under this setup, real-time markets are cleared independently for each $t\in\mathcal{T}$, and the expected revenues of solar panel investors are determined accordingly.
+In our extended model, each period $t$ is characterized by a distinct set of parameters: the total consumer load $L_{t}$, the solar generation per unit panel $G_{t}$ with associated probability density function $f_{G_{t}}$, and the utility supply price $\pi_{\mathrm{u}_{t}}$.
+In this model, real-time markets are cleared independently for each $t\in\mathcal{T}$, and the expected revenues of panel investors are determined accordingly.
 In contrast, contract-based markets involve long-term agreements, where consumers rent panels from suppliers once for the entire $\mathcal{T}$.
-Unlike the homogeneous case, where the expected revenue is identical across periods, here the revenues vary across operation periods.
-Hence, we compute them individually and then take their average.
-As in the homogeneous case, we then scale this average by $\widetilde{T}$ to obtain the expected total revenue $\Pi_{m}^{\mathrm{s}}(c)$.
+Unlike the homogeneous case, where expected revenue is identical across periods, revenues here vary by operation period.
+We therefore compute period-by-period revenues, average them, and then scale by $\widetilde{T}$ to obtain the expected total revenue $\Pi_{m}^{\mathrm{s}}(c)$.
 We start by characterizing the CE of short-term markets and associated revenue for investors under this new setup.
@@
 In the contract-based market, $d_{i}^{\star}(\pi)$ is a solution to
-$\max_{d_{i}\geq 0}\enspace\sum_{t\in\mathcal{T}}\left[v_{i}\mathbb{E}\min\{d_{i}G_{t},L_{t}\}-\pi_{\mathrm{u}_{t}}\mathbb{E}(L_{t}-d_{i}G_{t})_{+}\right]-\pi d_{i}.$
+$\max_{d_{i}\geq 0}\!\!\kern 5.0pt\sum_{t\in\mathcal{T}}\left[v_{i}\mathbb{E}\min\{d_{i}G_{t},L_{t}\}\!-\!\pi_{\mathrm{u}_{t}}\mathbb{E}(L_{t}-d_{i}G_{t})_{+}\right]\!-\!\pi d_{i}.$ (29)
 Similar to (11), for each individual $t\in\mathcal{T}$, we can define $\mu_{G_{t,\mathrm{tr}}}(d_{i}):=\mathbb{E}\left[G_{t}\mathbb{1}\bigl\{d_{i}G_{t}\leq L_{t}\bigl\}\right]$.
 Using this we can analytically characterize $d_{i}^{\star}(\pi)$ as a solution to $\sum_{t\in\mathcal{T}}\left(\pi_{\mathrm{u}_{t}}+v_{i}\right)\mu_{G_{t,\mathrm{tr}}}\left(d_{i}\right)=\pi.$ The left-hand side of this equation represents the consumer’s expected total valuation for one unit of rented panel capacity over the entire planning period.
-We treat this quantity as a monotonically decreasing function, denoted by $w_{i}(d_{i})$, and define an extension of its inverse as
+We treat this quantity as a non-increasing function, denoted by $w_{i}(d_{i})$, and define an extension of its inverse as
 […]
@@
 The social welfare optimal capacity, $c_{\mathrm{opt}}$, is a solution to
-$\max_{c\in\mathbb{R}_{+}}\enspace$ $\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\!\left[v_{t}^{\star}(c,G_{t})L_{t}\!-\!\pi_{\mathrm{u}_{t}}(L_{t}-cG_{t})_{+}\right]\!-\!\pi_{0}c.$ (28)
+$\max_{c\in\mathbb{R}_{+}}\kern 5.0pt$ $\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\!\left[v_{t}^{\star}(c,G_{t})L_{t}\!-\!\pi_{\mathrm{u}_{t}}(L_{t}-cG_{t})_{+}\right]\!-\!\pi_{0}c.$ (31)
 An analytical characterization of $c_{\mathrm{opt}}$ is then obtained by solving
@@
 VII-A Case Study and Problem Description
-We conduct numerical experiments in this section to assess our theoretical results.
-Given California’s leading role in solar adoption, we focus our analysis on this state and consider a large group of potential investors in late 2023 deciding whether to install solar panels across various locations within it.
-To simplify the analysis, we assume a single pooled distribution-level market for the entire state and consider the three representative market frameworks discussed earlier.
-We consider a 25-year planning horizon, with hourly operation periods.
-While our analytical results are applicable to any number of heterogeneous operation periods, we simplify the case study by grouping operation periods into two representative categories.
-The first group corresponds to daytime hours, between sunrise and sunset, when solar irradiation is available.
-The second group includes the remaining hours, during which solar irradiation is minimal or absent.
-This setup is equivalent to considering $\mathcal{T}=\{1,2\}$, where $t=1$ corresponds to the first group and $t=2$ to the second, with appropriate weights applied to each group in the expected revenue calculations.
-Solar irradiation data from 2000 to 2022 for San Francisco, California, are used to model the distributions $f_{G_{t}}(g)$ for all $t\in\mathcal{T}$ [25].
-To account for panels’ efficiency, all irradiation values are scaled by a factor of 0.2 to obtain the effective solar irradiation [26].
+In this section, we conduct numerical experiments to evaluate our theoretical results.
+Motivated by California’s leading solar adoption, we model a large population of potential investors in late 2023 deciding whether to install PV across locations statewide.
+To simplify the analysis, we assume a single statewide pooled distribution-level market and consider the three representative market frameworks introduced earlier.
+We consider a 25-year panel lifespan with hourly operation periods.
+Although our analysis allows for arbitrarily many heterogeneous periods, we group hours into two representative categories: daytime (sunrise-sunset) with solar irradiation, and nighttime with minimal or no irradiation.
+Equivalently, we set $\mathcal{T}=\{1,2\}$, where $t=1$ denotes daytime and $t=2$ nighttime, and weight each group in the expected-revenue calculations by its relative frequency (i.e., the number of hours it represents).
+We use solar irradiation data from 2000 to 2022 for San Francisco, California, to model $f_{G_{t}}(g)$ for all $t\in\mathcal{T}$ [24].
+To account for panel efficiency, we scale irradiation by a factor of $0.2$ to obtain the effective solar irradiation [25].
 A kernel density estimator is then used to fit a probability density function to the gathered data, modeling the daytime irradiation density $f_{G_{1}}(g)$.
@@
 Moreover, we set $L_{1}=27\,\mathrm{GWh}$ and $L_{2}=29\,\mathrm{GWh}$ [27].
-We consider an average capital and installation cost of \$2.7 per watt [30] of solar panels, which includes labor and equipment costs.
+We consider an average capital and installation cost of \$2.7 per watt [28] for solar panels, which includes labor and equipment costs.
 Figure 3: Probability density function for solar irradiation
@@
 VII-B1 Validity of Theorem 2
-Given this setup, we first compute the expected revenues of panel investors under the real-time markets using (25) and (26).
-Meanwhile, for the contract-based market, we obtain the aggregate demand function, i.e., $\widehat{d}^{\star}(\pi)$, using (27).
-Leveraging the characterizing equations for the heterogeneous case, we then obtain the NE aggregate investment capacity under the three considered short-term markets as summarized in Table II.
+Given this setup, we first compute panel investors’ expected revenues under the real-time markets using (27) and (28).
+For the contract-based market, we first obtain the aggregate demand function, $\widehat{d}^{\star}(\pi)$, from (30) and then compute investors’ revenue.
+Using the characterizing equations for the heterogeneous case, we derive the NE aggregate investment capacity under the three short-term markets, as summarized in Table II.
 TABLE II: NE aggregate investment capacity (GW)
@@
 We then consider the scenario where consumers assign no premium to solar energy, i.e., $\epsilon=0$, which yields identical aggregate capacities across all market mechanisms, as shown in Table II and implied by (26) under case 3 (no solar premium) of Theorems 1 and 2.
-In the following two subsections, we analyze how variations in consumers’ solar premiums and in panel capital and installation costs affect the comparison of equilibrium aggregate capacities across market mechanisms.
+As Table II suggests, discrepancies in equilibrium aggregate capacities are modest in the base case ($\epsilon=1$) because solar premiums are small on average (2.86 ¢\$/kWh) relative to the utility price (29 ¢\$/kWh).
+As premiums increase, these discrepancies are expected to widen, as analyzed next.
+In the following two subsections, we analyze how variations in consumers’ solar premiums and in panel capital and installation costs affect the equilibrium aggregate capacities.
 VII-B2 Sensitivity analysis with respect to consumers’ solar premiums
@@
 As consumers’ solar premiums increase, this figure reveals two key insights.
-First, the single-product real-time energy market results in a larger gap compared to the social welfare optimal capacity.
-This is primarily due to this market’s inability to capture the additional revenue potential from environmentally conscious consumers who are willing to pay a premium for clean energy.
+First, the single-product real-time market results in a larger gap compared to the social welfare optimal capacity.
+This is primarily due to this market’s inability to capture the additional revenue potential from environmentally conscious consumers.
 Second, the contract-based market consistently results in over-investment, reflected in the widening gap between $c^{\mathrm{ne}}_{\mathrm{prt}}$ and $c^{\mathrm{ne}}_{\mathrm{cb}}$.
-Figure 4: Variation of equilibrium solar panel investments in response to consumers’ solar premium scaling
+Figure 4: Variations of equilibrium aggregate panel investment capacities in response to consumers’ solar premium scaling
 VII-B3 Sensitivity analysis with respect to capital and installation costs
-We vary the capital and installation costs for a unit panel capacity and analyze how these changes affect the equilibrium aggregate capacities, as shown in Fig. 5.
-Figure 5: Equilibrium solar panel variation with changes in installation costs
+We vary the capital and installation costs for a unit panel capacity and analyze how these changes affect the equilibrium aggregate panel investment capacities, as shown in Fig. 5.
+Figure 5: Equilibrium aggregate panel investment capacities variation in response to changes in installation and capital costs
 An immediate interpretation of this figure is that increasing upfront costs reduces the aggregate installed capacity of solar panels, while maintaining the inequality $c_{\mathrm{srt}}^{\mathrm{ne}}<c_{\mathrm{opt}}=c_{\mathrm{prt}}^{\mathrm{ne}}<c_{\mathrm{cb}}^{\mathrm{ne}}$ up to a certain threshold.
@@
 VIII Concluding Remarks
-This paper examines how the design of short-term electricity markets influences long-term investment in distributed solar panels, where investors’ decisions are interdependent through the expected revenues earned from participating in these markets over the panels’ lifespans.
+This paper studies how short-term market mechanisms shape long-term distributed solar investment, where investors’ decisions are driven by expected lifetime revenue from participating in these markets.
 To analyze this relationship, we develop an investment model that links short-term market equilibria to the resulting long-term investment outcomes.
-As a key analytical tool, we introduce a unified modeling framework that derives closed-form expressions for short-term equilibrium outcomes across a broad class of market mechanisms.
-These expressions then serve as inputs to our investment model.
-Then, we focus on three representative short-term market mechanisms from the literature: a single-product real-time market, a product-differentiated real-time market, and a contract-based panel market, each reflecting a distinct approach to pricing and trading in distribution networks.
-For each, we apply our unified modeling framework followed by the investment model to derive analytical expressions for the corresponding long-term equilibrium panel capacity.
-We then theoretically establish that the product-differentiated real-time market leads to socially optimal investment, while the single-product market results in under-investment.
-Interestingly, when solar premiums of consumers are small, our theoretical analysis reveals that the contract-based market leads to over-investment, a pattern that our numerical results confirm even at higher levels of consumer premiums.
-Future directions include integrating centralized utility planning with the distributed investment problem and incorporating policy tools, such as tax incentives or feed-in tariffs, into our analysis of long-term investment equilibria.
-Appendix A Proofs for Section IV
+As a key tool, we propose a unified framework that yields closed-form expressions for short-term equilibrium outcomes across a broad class of mechanisms; these expressions feed into the investment model.
+We then analyze three representative mechanisms: a standard single-product real-time market, a new product-differentiated real-time market, and a new contract-based panel market, each capturing a distinct approach to pricing and trading in distribution networks.
+For each mechanism, we apply the unified framework and the investment model to derive analytical expressions for the resulting long-term equilibrium panel capacities.
+We theoretically establish that the product-differentiated real-time market yields socially optimal investment, whereas the single-product market induces under-investment.
+When consumers’ solar premiums are small, we further show that the contract-based market can lead to over-investment, a pattern our numerical results confirm even for larger premiums.
+Future work includes relaxing assumptions such as uniform buy/sell pricing in short-term markets, as well as incorporating investor heterogeneity (e.g., risk preferences).
+Appendix A Competitive Equilibrium and Nash Equilibrium Equivalence in Short-Term Markets
+As noted in Remark 2, rather than detailing the price-discovery process, we focus on the CE concept for the short-term markets, yet claim that it is equivalent to an NE of a bidding game between sellers and buyers.
+We establish this equivalence in this section.
+We consider a standard double-sided auction process for our unified solar market participation game.
+In this process, each seller $i\in\mathcal{I}_{\mathrm{s}}$ submits a price-quantity bid $(\underline{\pi}_{i}^{\mathrm{s}},\overline{q}^{\mathrm{s}}_{i})$, where $\underline{\pi}_{i}^{\mathrm{s}}\geq 0$ is the minimum price that the seller is willing to accept, and $\overline{q}^{\mathrm{s}}_{i}\in\mathcal{Q}^{\mathrm{s}}_{i}\subset\mathbb{R}_{+}$ is the maximum quantity that the seller can provide.
+Each buyer $i\in\mathcal{I}_{\mathrm{b}}$ submits a price-quantity offer $(\overline{\pi}_{i}^{\mathrm{b}},\overline{q}^{\mathrm{b}}_{i})$, where $\overline{\pi}_{i}^{\mathrm{b}}\geq 0$ is the maximum willingness to pay for buyer $i$, and $\overline{q}^{\mathrm{b}}_{i}\in\mathcal{Q}^{\mathrm{b}}_{i}\subset\mathbb{R}_{+}$ is the maximum quantity to be consumed by the buyer.
+We will refer to the collection of quantities for all sellers or buyers by omitting the subscript $i$, e.g., $\overline{q}^{\mathrm{s}}:=\{\overline{q}_{i}^{\mathrm{s}}\}_{i\in\mathcal{I}_{\mathrm{s}}}$.
+To further simplify our notation, we denote the collection of submitted market data by $\mathbf{D}=(\underline{\pi}^{\mathrm{s}},\overline{q}^{\mathrm{s}},\overline{\pi}^{\mathrm{b}},\overline{q}^{\mathrm{b}})$; we also use $\mathbf{D}_{i}$ to denote the data submitted by supplier $i\in\mathcal{I}_{\mathrm{s}}$ (or consumer $i\in\mathcal{I}_{\mathrm{b}}$), and $\mathbf{D}_{-i}$ to denote the data submitted by everyone else [^7] We will always make it clear that whether $i$ refers a supplier or consumer to avoid ambiguity in this notation.
+.
+We slightly overload notation and write the payoff of seller $i\in\mathcal{I}_{\mathrm{s}}$ as $\Pi_{i}^{\mathrm{s}}(\underline{\pi}_{i}^{\mathrm{s}},\overline{q}_{i}^{\mathrm{s}},\mathbf{D}_{-i})$, using the same payoff function $\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)$ defined in Section III for the corresponding market.
+Similarly, we denote the payoff of buyer $i\in\mathcal{I}_{\mathrm{b}}$ by $\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}},\overline{q}_{i}^{\mathrm{b}},\mathbf{D}_{-i})$, which is identical to the buyer payoff function $\Pi_{i}^{\mathrm{b}}(q_{i}^{\mathrm{b}},\pi)$ in Section III.
+The market operator will then collect all these bids and offers, and clear the market by identifying the intersection of the supply curve and demand curve obtained by integrating the bids and offers from individual participants.
+The market clearing process determines a market price $\pi$, and the allocation for all bids and offers, i.e., $q_{i}^{\mathrm{s}}\in[0,\overline{q}_{i}^{\mathrm{s}}]$ for all $i\in\mathcal{I}_{\mathrm{s}}$, and $q_{i}^{\mathrm{b}}\in[0,\overline{q}_{i}^{\mathrm{b}}]$ for all $i\in\mathcal{I}_{\mathrm{b}}$.
+In other words,
+$(\pi,q^{\mathrm{s}},q^{\mathrm{b}})=\mathsf{Market\,Clearing}(\mathbf{D}),$ (34)
+where the mathematical details for the market clearing process are provided as follows.
+A-A Market Clearing Process
+Given the bids and offers $\mathbf{D}=(\underline{\pi}^{\mathrm{s}},\overline{q}^{\mathrm{s}},\overline{\pi}^{\mathrm{b}},\overline{q}^{\mathrm{b}})$, the market operator clears the market by identifying the intersection of the supply and demand curves.
+It is easy to show that the following optimization is an equivalent formulation:
+$\max_{q^{\mathrm{s}},q^{\mathrm{b}}}\quad$ $\int_{\mathcal{I}_{\mathrm{b}}}\overline{\pi}^{\mathrm{b}}_{i}q_{i}^{\mathrm{b}}\,\,\mathrm{d}i-\int_{\mathcal{I}_{\mathrm{s}}}\underline{\pi}^{\mathrm{s}}_{i}q_{i}^{\mathrm{s}}\,\,\mathrm{d}i$ (35a)
+s.t. $\pi:\int_{\mathcal{I}_{\mathrm{s}}}q_{i}^{\mathrm{s}}\,\,\mathrm{d}i=\int_{\mathcal{I}_{\mathrm{b}}}q_{i}^{\mathrm{b}}\,\,\mathrm{d}i,$ (35b)
+   $\quad\,\,\,0\leq q^{\mathrm{s}}_{i}\leq\overline{q}^{\mathrm{s}}_{i},\quad i\in\mathcal{I}_{\mathrm{s}},$ (35c)
+   $\quad\,\,\,0\leq q^{\mathrm{b}}_{i}\leq\overline{q}^{\mathrm{b}}_{i},\quad i\in\mathcal{I}_{\mathrm{b}},$ (35d)
+where $\pi$ is the Lagrange multiplier associated with the market clearing constraint (35b).
+Optimization (35) is an infinite dimensional linear program, which has many solutions.
+To avoid the potential non-uniqueness issue in both the primal and dual solutions, we adapt the following rules:
+  1. 1.
+    If the dual solution of constraint (35b) is not unique, denote the set of such dual solutions by $\Omega^{\mathrm{eq}}$.
+We set the market price to $\pi:=\min\{\pi^{\prime}:\pi^{\prime}\in\Omega^{\mathrm{eq}}\}$.
+  2. 2.
+    For each buyer $i\in\mathcal{I}_{\mathrm{b}}$,
+    $q_{i}^{\mathrm{b}}=\begin{cases}\overline{q}^{\mathrm{b}}_{i},&\mbox{ if }\overline{\pi}^{\mathrm{b}}_{i}>\pi,\\ 0,&\mbox{ if }\overline{\pi}^{\mathrm{b}}_{i}<\pi.\end{cases}$
+    If both $i,j\in\mathcal{I}_{\mathrm{b}}$ such that $\overline{\pi}^{\mathrm{b}}_{i}=\overline{\pi}^{\mathrm{b}}_{j}=\pi$, then $q^{\mathrm{b}}_{i}/\overline{q}^{\mathrm{b}}_{i}=q^{\mathrm{b}}_{j}/\overline{q}^{\mathrm{b}}_{j}$.
+If $i\in\mathcal{I}_{\mathrm{b}}$ is the only buyer with $\overline{\pi}^{\mathrm{b}}_{i}=\pi$, then $q_{i}^{\mathrm{b}}=\overline{q}^{\mathrm{b}}_{i}$.
+  3. 3.
+    For each seller $i\in\mathcal{I}_{\mathrm{s}}$,
+    $q_{i}^{\mathrm{s}}=\begin{cases}\overline{q}^{\mathrm{s}}_{i},&\mbox{ if }\overline{\pi}^{\mathrm{s}}_{i}<\pi,\\ 0,&\mbox{ if }\overline{\pi}^{\mathrm{s}}_{i}>\pi.\end{cases}$
+    If both $i,j\in\mathcal{I}_{\mathrm{s}}$ such that $\overline{\pi}^{\mathrm{s}}_{i}=\overline{\pi}^{\mathrm{s}}_{j}=\pi$, then $q^{\mathrm{s}}_{i}/\overline{q}^{\mathrm{s}}_{i}=q^{\mathrm{s}}_{j}/\overline{q}^{\mathrm{s}}_{j}$.
+If $i\in\mathcal{I}_{\mathrm{s}}$ is the only seller with $\overline{\pi}^{\mathrm{s}}_{i}=\pi$, then $q_{i}^{\mathrm{s}}=\overline{q}^{\mathrm{s}}_{i}$.
+One can verify that the proposed rules above correspond to a solution of optimization (35).
+In the following two subsections, we characterize the Nash equilibrium of the buyer-seller bidding process under each of the three short-term markets, yielding a stable $\mathbf{D}$ for the market-clearing problem (35).
+A-B Nash Equilibrium of Real-Time Market Mechanisms and Its Equivalency to the Associated CE of Lemma 1
+Lemma 5 (NE of real-time market mechanisms)
+For product-differentiated real-time market (prt) and single-product real-time market (srt), an NE market participation strategy and the corresponding market clearing outcomes are listed in Table III.
+TABLE III: NE market participation strategy and market clearing outcomes for real-time market mechanisms
+NE participation strategy Market clearing outcome
+Market $\underline{\pi}_{i}^{\mathrm{s}}$ $\overline{q}_{i}^{\mathrm{s}}$ $\overline{\pi}_{i}^{\mathrm{b}}$ $\overline{q}_{i}^{\mathrm{b}}$ $\pi$ $q_{i}^{\mathrm{s}}$ $q_{i}^{\mathrm{b}}$
+Abundant supply $cG>L$ prt 0 cGcG $\pi_{\mathrm{u}}+v_{i}$ $L$ 0 $L$ $L$
+srt 0 cGcG $\pi_{\mathrm{u}}$ $L$ 0 $L$ $L$
+Limited supply $cG\leq L$ prt 0 cGcG $\pi_{\mathrm{u}}+v_{i}$ $L$ $\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)$ cGcG $L\mathbb{1}\{\overline{\pi}^{\mathrm{b}}_{i}\geq\pi\}$
+srt 0 cGcG $\pi_{\mathrm{u}}$ $L$ $\pi_{\mathrm{u}}$ cGcG cGcG
+Proof:
+We present the main proof for the product-differentiated real-time market; the single-product real-time market follows as a special case.
+We first show that, under the proposed participation strategies, the market outcomes in Table III solve (35).
+We then verify that these strategies constitute a Nash equilibrium, i.e., no supplier or consumer can profitably deviate unilaterally.
+Following Table III, we consider two cases.
+Abundant supply: The outcomes in Table III satisfy the constraints of (35) and are therefore feasible.
+To establish optimality, note that since $0\leq q_{i}^{\mathrm{b}}\leq L$ for all $i\in\mathcal{I}_{\mathrm{b}}$, the objective in (35) is bounded by
+$\int_{\mathcal{I}_{\mathrm{b}}}\left(\pi_{\mathrm{u}}+v_{i}\right)q_{i}^{\mathrm{b}}\,\mathrm{d}i\leq\int_{\mathcal{I}_{\mathrm{b}}}\left(\pi_{\mathrm{u}}+v_{i}\right)L\,\mathrm{d}i.$ (36)
+This upper bound is attained by the allocation $q_{i}^{\mathrm{b}}=L$ for all $i\in\mathcal{I}_{\mathrm{b}}$, hence the outcomes in Table III are optimal for (35).
+To derive the associated clearing price $\pi=0$ (as in Table III), we next apply the KKT conditions by forming the Lagrangian of (35) as
+$\mathcal{L}=$ $-\int_{\mathcal{I}_{\mathrm{b}}}\left(\pi_{\mathrm{u}}+v_{i}\right)q_{i}^{\mathrm{b}}\,\,\mathrm{d}i+\pi\left(\int_{\mathcal{I}_{\mathrm{b}}}q_{i}^{\mathrm{b}}\,\,\mathrm{d}i-\int_{\mathcal{I}_{\mathrm{s}}}q_{i}^{\mathrm{s}}\,\,\mathrm{d}i\right)$
+$+\int_{\mathcal{I}_{\mathrm{s}}}\mu_{i}^{\mathrm{s}}\left(q_{i}^{\mathrm{s}}-cG\right)\,\,\mathrm{d}i+\int_{\mathcal{I}_{\mathrm{b}}}\mu_{i}^{\mathrm{b}}\left(q_{i}^{\mathrm{b}}-L\right)\,\,\mathrm{d}i$
+$+\int_{\mathcal{I}_{\mathrm{s}}}\nu_{i}^{\mathrm{s}}\left(-q_{i}^{\mathrm{s}}\right)\,\,\mathrm{d}i+\int_{\mathcal{I}_{\mathrm{b}}}\nu_{i}^{\mathrm{b}}\left(-q_{i}^{\mathrm{b}}\right)\,\,\mathrm{d}i,$ (37)
+in which $\mu_{i}^{\mathrm{s}}$ and $\nu_{i}^{\mathrm{s}}$ are the Lagrangian multipliers associated with (35c), while $\mu_{i}^{\mathrm{b}}$ and $\nu_{i}^{\mathrm{b}}$ are for (35d).
+Differentiating $\mathcal{L}$ with respect to $q_{i}^{\mathrm{b}}$ and $q_{i}^{\mathrm{s}}$ yields
+$-\pi_{\mathrm{u}}-v_{i}+\pi+\mu_{i}^{\mathrm{b}}-\nu_{i}^{\mathrm{b}}=0,$ $i\in\mathcal{I}_{\mathrm{b}},$
+$-\pi+\mu_{i}^{\mathrm{s}}-\nu_{i}^{\mathrm{s}}=0,$ $i\in\mathcal{I}_{\mathrm{s}}.$
+Given the clearing outcomes in Table III, the constraints associated with $\mu_{i}^{\mathrm{s}}$, $\nu_{i}^{\mathrm{s}}$, and $\nu_{i}^{\mathrm{b}}$ are nonbinding; hence, by complementary slackness, $\mu_{i}^{\mathrm{s}}=\nu_{i}^{\mathrm{s}}=\nu_{i}^{\mathrm{b}}=0$.
+Therefore,
+$\pi=\pi_{\mathrm{u}}+v_{i}-\mu_{i}^{\mathrm{b}},$ $i\in\mathcal{I}_{\mathrm{b}},$
+$\pi=0,$ $i\in\mathcal{I}_{\mathrm{s}},$
+which implies $\pi=0$.
+Thus, the clearing outcomes in Table III are optimal for (35) and therefore constitute the market-clearing solution in this case.
+We next show that the corresponding submitted bids form a Nash equilibrium of the bidding process.
+Under the results summarized for this case in Table III, each supplier earns $\Pi_{i}^{\mathrm{s}}(0,cG,\mathbf{D}_{-i})=0$.
+Suppose, for contradiction, that some $i\in\mathcal{I}_{\mathrm{s}}$ deviates to $\underline{\pi}_{i}^{\mathrm{s}^{\prime}}\neq 0$ and obtains $\Pi_{i}^{\mathrm{s}}(\underline{\pi}_{i}^{\mathrm{s}^{\prime}},cG,\mathbf{D}_{-i})>0$, which would require a strictly positive clearing price $\pi^{\prime}>0$.
+However, with all other bids fixed, a single seller’s deviation does not affect the aggregate terms in (35a)-(35b); hence the corresponding KKT multiplier of (35b) (the market price) remains unchanged.
+Therefore $\pi^{\prime}=\pi=0$ and the deviating seller’s payoff stays zero, a contradiction.
+Similarly, suppose that there exists an $i\in\mathcal{I}_{\mathrm{b}}$ with $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}\neq\pi_{\mathrm{u}}$, and $\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},L,\mathbf{D}_{-i})>\Pi_{i}^{\mathrm{b}}(\pi_{\mathrm{u}}+v_{i},L,\mathbf{D}_{-i}).$
+However, using (5) gives
+$\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},L,\mathbf{D}_{-i})\leq v_{i}q_{i}^{\mathrm{b}},$
+and
+$\Pi_{i}^{\mathrm{b}}(\pi_{\mathrm{u}}+v_{i},L,\mathbf{D}_{-i})=v_{i}q_{i}^{\mathrm{b}}.$
+Therefore, $\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},L,\mathbf{D}_{-i})\leq\Pi_{i}^{\mathrm{b}}(\pi_{\mathrm{u}}+v_{i},L,\mathbf{D}_{-i})$ which contradicts with the initial assumption.
+Thus, no supplier or consumer has a profitable unilateral deviation, and the outcomes in Table III constitute a Nash equilibrium for this case.
+The single-product real-time market follows by setting $v_{i}\equiv 0$.
+Limited supply: To verify that the market outcomes in Table III clear the market under the proposed bids, it suffices to show that they satisfy the KKT conditions of (35).
+We therefore form the Lagrangian as in (A-B) and obtain the stationarity conditions as follows:
+$v_{i}-\overline{F}_{V}^{-1}\!\left(\frac{cG}{L}\right)=\mu_{i}^{\mathrm{b}}-\nu_{i}^{\mathrm{b}}\triangleq\Upsilon_{i}^{\mathrm{b}},$ $i\in\mathcal{I}_{\mathrm{b}},$
+$\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\!\left(\frac{cG}{L}\right)=\mu_{i}^{\mathrm{s}}-\nu_{i}^{\mathrm{s}},$ $i\in\mathcal{I}_{\mathrm{s}},$ (38)
+where $\Upsilon_{i}^{\mathrm{b}}\in\mathbb{R}$ is introduced for notational convenience. Moreover, the given bids also satisfy the primal feasibility conditions:
+$0\leq L\mathbb{1}\left\{v_{i}\geq\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right\}\leq L,$
+$0\leq cG\leq cG,$
+$cG=\int_{\mathcal{I}_{\mathrm{b}}}L\mathbb{1}\left\{v_{i}\geq\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right\}\,\,\mathrm{d}i.$
+The complementary slackness conditions associated with (35c) can be written as:
+$\nu_{i}^{\mathrm{s}}\bigl(-cG\bigl)=0\Rightarrow\nu_{i}^{\mathrm{s}}=0,$ $\quad i\in\mathcal{I}_{\mathrm{s}},$
+$\mu_{i}^{\mathrm{s}}\bigl(cG-cG\bigl)=0\Rightarrow\mu_{i}^{\mathrm{s}}\geq 0,$ $\quad i\in\mathcal{I}_{\mathrm{s}}.$
+Based on these results and according to (38), it is immediate to get
+$\mu_{i}^{\mathrm{s}}=\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right),\quad i\in\mathcal{I}_{\mathrm{s}}.$
+Similarly, the complementary slackness conditions associated with (35d) can be written as:
+$\nu_{i}^{\mathrm{b}}\left(-L\mathbb{1}\left\{v_{i}\geq\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right\}\right)=0,$ $\quad i\in\mathcal{I}_{\mathrm{b}},$
+$\mu_{i}^{\mathrm{b}}\left(L\mathbb{1}\left\{v_{i}\geq\overline{F}_{V}^{-1}\left(\frac{cG}{L}\right)\right\}-L\right)=0,$ $\quad i\in\mathcal{I}_{\mathrm{b}}.$
+In particular, letting $\mu_{i}^{\mathrm{b}}=(\Upsilon_{i}^{\mathrm{b}})_{+}$ and $\nu_{i}^{\mathrm{b}}=(-\Upsilon_{i}^{\mathrm{b}})_{+}$ satisfies dual feasibility and complementary slackness.
+Therefore, under the proposed bids, the claimed market outcomes satisfy the KKT conditions and hence are feasible and optimal solutions of (35).
+To establish that the proposed strategies form a Nash equilibrium, suppose for contradiction that some $i\in\mathcal{I}_{\mathrm{s}}$ deviates to $\underline{\pi}_{i}^{\mathrm{s}^{\prime}}\neq 0$ and attains
+$\Pi_{i}^{\mathrm{s}}(\underline{\pi}_{i}^{\mathrm{s}^{\prime}},cG,\mathbf{D}_{-i})>\Pi_{i}^{\mathrm{s}}(0,cG,\mathbf{D}_{-i})=\Bigl(\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\!\Bigl(\tfrac{cG}{L}\Bigr)\Bigr)cG.$
+Such a profitable deviation would require the clearing price to rise to some $\pi^{\prime}>\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\!\bigl(\tfrac{cG}{L}\bigr)$, which is impossible since any individual participant has a negligible impact on the market price, as discussed in the previous case.
+Similarly, suppose for contradiction that there exists $i\in\mathcal{I}_{\mathrm{b}}$ with $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}\neq\pi_{\mathrm{u}}+v_{i}$ and $\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},L,\mathbf{D}_{-i})>\Pi_{i}^{\mathrm{b}}(\pi_{\mathrm{u}}+v_{i},L,\mathbf{D}_{-i})$.
+We distinguish two cases:
+  1. 1.
+    If $\pi_{i}^{\mathrm{b}}\geq\pi$, then
+    $\Pi_{i}^{\mathrm{b}}(\pi_{\mathrm{u}}+v_{i},L,\mathbf{D}_{-i})=\Bigl(v_{i}-\pi_{\mathrm{u}}-\overline{F}_{V}^{-1}\!\Bigl(\tfrac{cG}{L}\Bigr)\Bigr)L.$
+    If $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}\geq\pi_{i}^{\mathrm{b}}$, the deviation does not affect the clearing price (as argued before), hence yields the same payoff:
+    $\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},L,\mathbf{D}_{-i})=\Bigl(v_{i}-\pi_{\mathrm{u}}-\overline{F}_{V}^{-1}\!\Bigl(\tfrac{cG}{L}\Bigr)\Bigr)L,$
+    contradicting strict improvement.
+The same conclusion holds for $\pi<\overline{\pi}_{i}^{\mathrm{b}^{\prime}}<\pi_{i}^{\mathrm{b}}$.
+If instead $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}<\pi$, then $\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},L,\mathbf{D}_{-i})=-\pi_{\mathrm{u}}L$, again contradicting the assumed profitable deviation.
+  2. 2.
+    If $\pi_{i}^{\mathrm{b}}<\pi$, then
+    $\Pi_{i}^{\mathrm{b}}(\pi_{\mathrm{u}}+v_{i},L,\mathbf{D}_{-i})=-\pi_{\mathrm{u}}L.$
+    For any $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}<\pi$ (including $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}<\pi_{i}^{\mathrm{b}}$ or $\pi_{i}^{\mathrm{b}}<\overline{\pi}_{i}^{\mathrm{b}^{\prime}}<\pi$), the payoff remains $\pi_{\mathrm{u}}L$, contradicting strict improvement.
+If $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}>\pi$, then
+    $\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},L,\mathbf{D}_{-i})=\Bigl(v_{i}-\pi_{\mathrm{u}}-\overline{F}_{V}^{-1}\!\Bigl(\tfrac{cG}{L}\Bigr)\Bigr)L.$
+    For this to exceed $\pi_{\mathrm{u}}L$, it must hold that
+    $\Bigl(v_{i}-\pi_{\mathrm{u}}-\overline{F}_{V}^{-1}\!\Bigl(\tfrac{cG}{L}\Bigr)\Bigr)L+\pi_{\mathrm{u}}L>0,$ (39)
+    which implies $v_{i}-\overline{F}_{V}^{-1}\!\bigl(\tfrac{cG}{L}\bigr)>0$. However, since we are in the case that $\pi_{i}^{\mathrm{b}}<\pi$, it gives
+    $\pi_{\mathrm{u}}+v_{i}-\pi=\pi_{\mathrm{u}}+v_{i}-\Bigl(\pi_{\mathrm{u}}+\overline{F}_{V}^{-1}\!\Bigl(\tfrac{cG}{L}\Bigr)\Bigr)<0,$
+    and hence $v_{i}-\overline{F}_{V}^{-1}(\tfrac{cG}{L})<0$, contradicting (39).
+Therefore, no buyer or seller admits a profitable unilateral deviation, and the strategies in Table III form a Nash equilibrium in this case.
+The single-product real-time market follows by setting $v_{i}\equiv 0$.
+Having shown that the bidding outcomes in Table III constitute an Nash equilibrium and clear the market, we compare them with the CE outcomes in Table I.
+Since the two coincide, the Nash equilibrium of the bidding process is equivalent to the CE characterized in Lemma 1 for the real-time mechanisms.
+We next establish the same equivalence for the contract-based market.
+A-C Nash Equilibrium of Contract-based Market Mechanism and Its Equivalency to the Associated CE of Lemma 3
+Lemma 6 (NE for the contract-based market)
+The following states a Nash equilibrium for the contract-based market and the corresponding clearing outcomes:
+$(\underline{\pi}_{i}^{\mathrm{s}},\overline{q}_{i}^{\mathrm{s}})=(0,c),$ $q_{i}^{\mathrm{s}}=c,\qquad$ $\qquad i\in\mathcal{I}_{\mathrm{s}},$
+$(\overline{\pi}_{i}^{\mathrm{b}},\overline{q}_{i}^{\mathrm{b}})=(\pi,d_{i}^{\star}(\pi)),$ $q_{i}^{\mathrm{b}}=d_{i}^{\star}(\pi),$ $\qquad i\in\mathcal{I}_{\mathrm{b}},$
+and $\pi$ is the solution to $c=\widehat{d}^{\star}(\pi).$
+Proof:
+We first verify that, under the proposed strategies, the claimed market outcome is valid by showing it solves (35).
+To this end, note that the outcome clearly satisfies all constraints of (35) and is therefore feasible.
+Moreover, under the assumed strategies, the objective in (35) is bounded by
+$\int_{\mathcal{I}_{\mathrm{b}}}\pi q_{i}^{\mathrm{b}}\,\mathrm{d}i\leq\int_{\mathcal{I}_{\mathrm{b}}}\pi d_{i}^{\star}(\pi)\,\mathrm{d}i.$
+Since the claimed outcome attains this bound, it is optimal for (35).
+To establish that these strategies form a Nash equilibrium, we first show that no supplier has a profitable unilateral deviation, using the same argument as in the real-time market proof.
+For consumers, note that for any given price $\pi$, each buyer $i\in\mathcal{I}_{\mathrm{b}}$ maximizes its payoff by choosing $d_{i}^{\star}(\pi)$, the optimizer of (7).
+Hence, there is no $\overline{q}_{i}^{\mathrm{b}^{\prime}}\neq d_{i}^{\star}(\pi)$ such that $\Pi_{i}^{\mathrm{b}}(\pi,\overline{q}_{i}^{\mathrm{b}^{\prime}},\mathbf{D}_{-i})>\Pi_{i}^{\mathrm{b}}(\pi,d_{i}^{\star}(\pi),\mathbf{D}_{-i})$.
+Given this, it remains to show that the equilibrium bid satisfies $\overline{\pi}_{i}^{\mathrm{b}}=\pi$.
+Suppose, for contradiction, that some $i\in\mathcal{I}_{\mathrm{b}}$ submits $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}\neq\pi$.
+The bidder then chooses the corresponding quantity $\overline{q}_{i}^{\mathrm{b}^{\prime}}=d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}})$, i.e., the optimizer of (7) evaluated at its perceived price.
+[^8] Although an infinitesimal buyer cannot affect the clearing price, its price bid determines the quantity it offers.
+Assume further that this deviation is profitable:
+$\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}}),\mathbf{D}_{-i})>\Pi_{i}^{\mathrm{b}}(\pi,d_{i}^{\star}(\pi),\mathbf{D}_{-i}).$ (40)
+If $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}<\pi$, then by the market rule $q_{i}^{\mathrm{b}}=0$, so the deviation cannot improve payoff.
+If instead $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}>\pi$, the buyer is cleared at $q_{i}^{\mathrm{b}^{\prime}}=d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}})$.
+Since $d_{i}^{\star}(\cdot)$ is decreasing by (12) and $\overline{\pi}_{i}^{\mathrm{b}^{\prime}}>\pi$, we get $d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}})\leq d_{i}^{\star}(\pi)$, and hence
+$v_{i}\mathbb{E}\min\{d_{i}^{\star}(\pi)G,L\}-\pi\,d_{i}^{\star}(\pi)-\pi_{\mathrm{u}}\mathbb{E}\bigl(L-d_{i}^{\star}(\pi)G\bigr)_{+}>$
+$v_{i}\mathbb{E}\min\{d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}})G,L\}-\pi\,d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}})-\pi_{\mathrm{u}}\mathbb{E}\bigl(L-d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}})G\bigr)_{+}.$ (41)
+By (7), the left- and right-hand sides of (A-C) are $\Pi_{i}^{\mathrm{b}}(\pi,d_{i}^{\star}(\pi),\mathbf{D}_{-i})$ and $\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}}),\mathbf{D}_{-i})$, respectively, implying
+$\Pi_{i}^{\mathrm{b}}(\pi,d_{i}^{\star}(\pi),\mathbf{D}_{-i})>\Pi_{i}^{\mathrm{b}}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}},d_{i}^{\star}(\overline{\pi}_{i}^{\mathrm{b}^{\prime}}),\mathbf{D}_{-i}),$ (42)
+which contradicts (40). Thus, no buyer or seller has a profitable unilateral deviation from the strategies given in Lemma 6, and they therefore constitute a Nash equilibrium. ∎
+Comparing the market outcomes in Lemma 6 with the CE characterization in Lemma 3 shows that the proposed Nash equilibrium coincides with the CE for the contract-based market.
+Appendix B Proofs for Section IV
 Proof:
@@
 These results yield $\Pi_{i}^{\mathrm{s}}(q_{i}^{\mathrm{s}},\pi)$ for all $i\in\mathcal{I}_{\mathrm{s}}$, which can then be used in (8) to derive (14).
-Appendix B Proofs for section V
-B-A Proof of Lemma 4 and Proposition 1
+Appendix C Proofs for section V
+C-A Proof of Lemma 4 and Proposition 1
 Proof:
@@
 which completes the proof. ∎
-B-B Proof of Theorem 1
+C-B Proof of Theorem 1
 We establish each case individually:
@@
 The main steps of the proof are summarized as follows: 1) we treat the left-hand sides of (44) and (47) as functions of both $\epsilon$ (via its influence on $v$) and $c$; 2) we show that if $\epsilon$ lies in a neighborhood of 0, then the corresponding capacities (i.e., the values of $c$ that solve (44) and (47)) must lie in a neighborhood of $c_{0}$; 3) we derive the first-order Taylor expansions of these functions at $(\epsilon,c)=(0,c_{0})$, and provide explicit bounds on the higher-order terms; and 4) we compare the resulting Taylor expansions term by term.
-We defer the detailed proofs of the first three steps to Appendix C, and summarize their results below in the form of lemmas.
+We defer the detailed proofs of the first three steps to Appendix D, and summarize their results below in the form of lemmas.
 Lemma 7 (First-order expansion of $c_{\mathrm{prt}}^{\mathrm{ne}}$ around $\epsilon=0$)
@@
 $\beta:=\frac{\left(1-\lambda\right)\mathbb{E}\left[\overline{F}_{\widetilde{V}}^{-1}\left(\frac{c_{0}G}{L}\right)\,G\,\mathbb{1}\bigl\{c_{0}G\leq L\bigl\}\right]}{-\left(1+\lambda\right)\pi_{\mathrm{u}}\mu_{G_{\mathrm{tr}}}^{\prime}(c_{0})}>0.$
-Appendix C Proof of Lemma 5 and Lemma 6
-C-A Proof of Lemma 5
+Appendix D Proof of Lemma 7 and Lemma 8
+D-A Proof of Lemma 7
 We can use (11) to reformulate (44) as
@@
 $\lVert(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})-(0,c_{0})\rVert^{2}\leq\epsilon^{2}\left(1+L_{\Psi}^{2}\right),$
-which can be applied in (C-A) to reformulate it as
+which can be applied in (D-A) to reformulate it as
 $\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}})=$ $-\epsilon\,k\left(c_{0}\right)-\pi_{\mathrm{u}}\,(c^{\mathrm{ne}}_{\mathrm{prt}}-c_{0})\,\mu_{G_{\mathrm{tr}}}^{\prime}(c_{0})$
@@
 $0<m_{2}\leq\epsilon\left(\frac{k(c_{2})-k(c_{1})}{c_{1}-c_{2}}\right)\leq M_{2}.$ (67)
-Adding inequalities (53) and (55), and letting $m=m_{1}+m_{2}$ and $M=M_{1}+M_{2}$, we obtain (C-A), which in turn implies (50).
+Adding inequalities (65) and (67), and letting $m=m_{1}+m_{2}$ and $M=M_{1}+M_{2}$, we obtain (D-A), which in turn implies (62).
 Next, observe that $\phi_{\mathrm{prt}}(\epsilon,c^{\mathrm{ne}}_{\mathrm{prt}}):[0,1]\times\mathbb{R}\to\mathbb{R}$ is continuous, and $[0,1]$ is a compact set.
@@
 which completes the proof of Lemma 9.
-C-B Proof of Lemma 6
+D-B Proof of Lemma 8
 Using (11), we can reconstruct (47) as:
@@
 $\left\lVert\left(\epsilon,d_{i}^{\star}\left(\frac{\pi_{0}}{\widetilde{T}}\right)\right)-(0,c_{0})\right\rVert^{2}\leq\epsilon^{2}\left(1+\left(L_{\mathrm{cb},i}\right)^{2}\right).$
-Using this result in (C-B), we can reformulate (C-B) as
+Using this result in (D-B), we can reformulate (D-B) as
 […]
@@
 […]
-Appendix D Proofs for Section VI
-D-A Proof of Proposition 2
+Appendix E Proofs for Section VI
+E-A Proof of Proposition 2
 The expression for $v_{t}^{\star}(c,G_{t})$ is derived individually for each $t\in\mathcal{T}$ using the same method as in the proof of Proposition 1. Substituting these expressions into (31) yields:
-$\max_{c\in\mathbb{R}_{+}}\enspace\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\Biggr{[}$ $L_{t}\mathbb{1}\!\left\{cG_{t}\leq L_{t}\right\}\!\int_{0}^{F_{V}^{-1}\left(1-\frac{cG_{t}}{L_{t}}\right)}v_{i}\,\,\mathrm{d}F_{V}(v_{i})$
+$\max_{c\in\mathbb{R}_{+}}\kern 5.0pt\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\Biggr[$ $L_{t}\mathbb{1}\!\left\{cG_{t}\leq L_{t}\right\}\!\int_{0}^{F_{V}^{-1}\left(1-\frac{cG_{t}}{L_{t}}\right)}v_{i}\,\,\mathrm{d}F_{V}(v_{i})$
 $+L_{t}\mathbb{E}[V]-\pi_{\mathrm{u_{t}}}\mathbb{(}L_{t}-cG_{t})_{+}\Biggr]-\pi_{0}c.$
@@
 $\blacksquare$
-D-B Proof of Theorem 2
+E-B Proof of Theorem 2
 We prove each case individually:
@@
 Proof:
-The steps follow the same structure as in the proof of Theorem 1. Therefore, we begin by introducing new lemmas analogous to Lemmas 5 and 6.
+The steps follow the same structure as in the proof of Theorem 1. Therefore, we begin by introducing new lemmas analogous to Lemmas 7 and 8.
 Lemma 10 (First-order approximation of $c_{\mathrm{prt}}^{\mathrm{ne}}$ around $\epsilon=0$ with heterogeneous periods )
@@
 We then apply the result from (51) to each operation period; summing the resulting inequalities yields the above result and completes the proof. ∎
-Appendix E Proof of Lemma 8 and Lemma 9
-E-A Proof of Lemma 8
+Appendix F Proof of Lemma 10 and Lemma 11
+F-A Proof of Lemma 10
 We can rewrite (74) as:
@@
 […]
-E-B Proof of Lemma 9
+F-B Proof of Lemma 11
 We begin by rewriting (80) in the following form: